### v1

# Learning “Partner-Aware” Collaborators in Multi-Party Collaboration

Abstract
Large Language Models (LLMs) are increasingly bring deployed in agentic settings where they act as collaborators with humans. Therefore, it is increasingly important to be able to evaluate their abilities to collaborate effectively in multi-turn, multi-party tasks. In this paper, we build on the AI alignment and “safe interruptability” literature to offer novel theoretical insights on collaborative behavior between LLM-driven collaborator agents and an intervention agent. Our goal is to learn an ideal “partner-aware” collaborator that increases the group’s common-ground (CG)—alignment on task-relevant propositions—by intelligently collecting information provided in interventions by a partner agent.We show how LLM agents trained using standard RLHF and related approaches are naturally inclined to ignore possibly well-meaning interventions, which makes increasing group common ground non-trivial in this setting. We employ a two-player Modified-Action MDP to examine this suboptimal behavior of standard AI agents, and propose Interruptible Collaborative Roleplayer (ICR)—a novel “partner-aware” learning algorithm to train CG-optimal collaborators. Experiments on multiple collaborative task environments show that ICR, on average, is more capable of promoting successful CG convergence and exploring more diverse solutions in such tasks.
1 Introduction
As Large Language Models (LLMs) become rapidly integrated into workflows in various domains, such as educational settings and the workplace [Xiao et al., 2023], they are increasingly being deployed as “agents” that collaborate with humans using both general-purpose assistance [Grassucci et al., 2025] and task-specific support [Alhafni et al., 2024]. In these settings they often adopt “roles” or personalities [Li et al., 2023, Tseng et al., 2024, Hao et al., 2024, Kim et al., 2024] that can be flexibly assigned by human users.
Small-group collaborative settings (e.g., Karadzhov et al. [2023], Khebour et al. [2024a]), present unique opportunities for studying intelligent agent behavior in cooperative environments where participants deliberate to reconcile different assumptions and beliefs. During such collaborations, participants naturally encounter reasoning challenges stemming from task complexity, communication ambiguities, or cognitive biases. In these scenarios, interventions—suggestions or clarifications from collaborative agents—can significantly enhance task success by promoting “slow thinking” [Kahneman, 2011] and promoting the growth of common ground [Stalnaker, 2002]. Consider, for example, a group of students collaborating in a classroom science lab to determine the volume of an object by the amount of water it displaces. An assistive AI agent or more experienced peer might intervene with suggestions to help scaffold collaborative reasoning. However, poorly-timed interventions may interrupt collaborative flow, and misleading interventions can be detrimental [Peters et al., 2017a]. As learners, the students have incomplete knowledge, and so they may make their own suggestions under incorrect assumptions, or they may interpret their partners’ suggestions through the lens of their current presuppositions (for example, assuming that heavier objects must be more dense). This creates a fundamental challenge: how can we develop collaborator agents that effectively distinguish between helpful interventions and those that are poorly-grounded, based on flawed reasoning, or uncritically incorporating irrelevant or misleading context? A successful partner-aware collaborator agent would be able to include its understanding of its interlocutors’ beliefs to accurately interpret what in its partner’s suggestions can be taken at face value to steer their understanding toward learning gains based on what they already know, and what parts of an intervention or suggestion may be misleading or deepen misunderstanding. In this work, we address this critical question by developing a principled approach to train counterfactually-robust AI collaborators—agents that maintain logical consistency and task focus despite potentially misleading interventions from other participants.
We hypothesize that optimizing for general task utility (e.g., interventions that ultimately lead to correct task solutions) through counterfactual regularization encourages “partner-aware” behavior, leading to higher common ground convergence. Importantly, under our hypothesis, a true collaborator agent itself never has any more information than the aggregate of the group, and so common ground convergence should occur even without explicitly training for it. That is, an intentional collaborator learns to adapt: integrating helpful interventions while critically evaluating flawed ones. This ability to distinguish signal from noise fosters belief alignment as an emergent property of training, with practical benefits. In zero-shot or real-world collaborative settings, where intervention styles or partners are unfamiliar, counterfactually-trained agents should generalize better by leveraging learned notions of intervention quality. We validate this through a method we call Interruptible Collaborative Roleplayer (ICR), where we withhold common ground-based rewards during training and showing that such agents still achieve greater convergence than sophisticated LLM-agent training baselines, suggesting they have internalized collaboration principles transferable across partners and task to “in-the-wild” settings. Our work advances the state of the art in LLM-based collaborative agents through the following contributions [^1] Our code is available at https://github.com/csu-signal/ICR:
  • A novel theoretical framework that combines (1) a Modified-Action MDP (MAMDP) formulation explicitly modeling collaborator-intervention dynamics at the utterance or intervention level, and (2) a principled counterfactual invariance objective that regularizes the collaborator’s policy to remain consistent even when the specific influence pathway [Farquhar et al., 2022] of an intervention is nullified, via a simple counterfactual prompt prefix. Unlike prior approaches to multi-agent interaction [Langlois and Everitt, 2021, Jaques et al., 2019], our formulation specifically addresses the challenge of maintaining robust reasoning in the face of potentially misleading interventions.
  • Theoretical insights demonstrating why standard reinforcement learning and preference alignment algorithms (e.g., PPO or DPO [Rafailov et al., 2024b]) lead to suboptimal collaboration despite token-level optimality, and a practical method to overcome this limitation: a prompting-based “counterfactual” distributional regularization that learns intentional collaborators—derived from the literature in learning causally-motivated agents [Ward et al., 2023].
  • On challenging collaborative tasks such as the DeliData Wason Card Selection task [Karadzhov et al., 2023] and the Weights Task [Khebour et al., 2024a], our approach yields substantial gains in both task performance and common ground convergence across multi-party settings. Crucially, these improvements hold across both language-rich (full-press) and language-free (no-press) conditions, demonstrating the robustness of our collaborator agents. Our collaborator agents effectively distinguish between helpful and misleading interventions, maintaining logical consistency while benefiting from truly valuable input.
2 Related Work
Collaborative Reasoning and Interruptibility
While interruptibility has been studied in safety-critical RL [Orseau and Armstrong, 2016, Hadfield-Menell et al., 2017], it is equally vital in collaborative dialogue, where agents must discern whether interventions aid or hinder shared understanding [Grice, 1975, Sutton and Rao, 2024]. Prior work has explored these ideas in adversarial or game-theoretic contexts [Langlois and Everitt, 2021, Ward et al., 2023], but less so in multi-party deliberative language settings [Nath et al., 2025a, Obiso et al., 2025]. We extend this by training collaborator agents that are counterfactually robust—they update their beliefs when interventions are helpful, while resisting misleading or misaligned input.
Text-based Agents and Collaborative Games
Text-grounded agents have been studied extensively in tool-use [Schick et al., 2023, Yao et al., 2022], navigation [Zhou et al., 2023], programming [Yang et al., 2023, Li et al., 2022, Lin et al., 2018], and roleplay [Li et al., 2023, Tseng et al., 2024], including multi-agent settings [Jiang et al., 2024]. While much of this focuses on single-agent optimization, collaborative games—such as Diplomacy [FAIR et al., 2022] and the Wason Card Selection task in the DeliData dataset [Karadzhov et al., 2023]—involve language-mediated belief alignment. In these domains, interruptions are rare [Peters et al., 2017b, Puranik et al., 2020], yet critical for resolving misunderstandings. More importantly, real-world datasets are textually sparse [Khebour et al., 2024a] or lack diversity in failure examples [Nokes-Malach et al., 2012]. Our work addresses this by providing a principled simulation-based method to collect two-way “expert”-AI interactions, which our ICR method stays integrated with at test-time.
Preference Learning and LLM Alignment
Preference-based LLM alignment with human intent [Christiano et al., 2017, Ziegler et al., 2020, Casper et al., 2023] has seen more efficient offline variants such as DPO [Rafailov et al., 2024b], IPO [Azar et al., 2024], and ORPO [Hong et al., 2024] that extend this by optimizing over contrastive pairs, avoiding the instability of full RL [Schulman et al., 2017]. These have been applied to many language tasks [Xu et al., 2024, Wei et al., 2023, Chen et al., 2024, Choi et al., 2024, Zhang et al., 2024], but little work targets multi-agent collaborative reasoning. Unlike information-seeking agents [Abdulhai et al., 2023, Andukuri et al., 2024], good collaborators must balance accuracy and consensus-building, especially over multiple interaction turns. Recent work [Rafailov et al., 2024a, Song et al., 2024] provide insights into how methods like DPO can be seen as “token-MDPs” that model multi-turn interactions [Sutton and Barto, 2018, Zhou et al., 2024] and likely do credit assignment. This relates to causal and counterfactual methods [Pearl, 2009, Ward et al., 2023, Wang et al., 2025] that test for beliefs, desires and intentions (BDI) [Bratman, 1987, Halpern and Kleiman-Weiner, 2018] in LLM-agents, and assign “intention” to parametric agents using Path-Specific Objectives (PSO) [Farquhar et al., 2022]. We extend this line of work with a principled yet efficient way for collaborator agents to explicitly regularize against a counterfactual policy, addressing limitations that emerge when collaborations are paired with an autonomous intervention agent and are required to be optimal over the space of interventional utterances.
3 The Collaborator’s Dilemma
Training LLMs to act as robust multiparty collaborator agents poses several fundamental challenges. First, high-quality human data on collaborative decision-making is limited, which restricts the scalability of supervised approaches for LLMs [Shih et al., 2021] using human-prior based learning techniques like InstructRL [Hu and Sadigh, 2023]. Secondly, successful collaborator agents must exhibit generalizability—they need to adapt to the diverse styles and conventions of their partners (both fellow task-focused collaborators and distinct intervention agents) to foster effective coordination. This adaptability should allow them to leverage prior experiences with similar partners on new tasks, while also retaining core task-specific skills when paired with entirely new partners.
At the heart of this challenge lies a key intuition: collaborators should not naively follow interventions exactly as intended by the intervening agent [Orseau and Armstrong, 2016, Hadfield-Menell et al., 2017]. In realistic dialogue settings, collaborators often reinterpret, resist, or transform interventions [Grice, 1975] in light of their internal goals—a process akin to belief revision [Bolander, 2014]. Robust collaboration requires identifying and incorporating helpful interventions, while critically evaluating or discarding those that are misaligned, manipulative, or simply incorrect (e.g., LLM hallucinations). However, this discernment is difficult because the collaborator typically lacks access to the intervener’s internal reward function or reliability about the intervener’s ultimate goal/objective.
To capture this interactional asymmetry, we adopt the Modified-Action Markov Decision Process (MAMDP) framework [Langlois and Everitt, 2021, Everitt et al., 2021] [^2] While two-player Markov Games are standard in MARL [Hu and Sadigh, 2023], the MAMDP offers a more intuitive fit for autoregressive LLMs by allowing the intervention policy $\pi_{I}$ to be fixed., modeling the interaction between a trained collaborator agent $\pi_{C}$ and an intervention agent $\pi_{I}$ as $M=(S,A_{C},A_{I},P_{S},P_{A},R,\gamma)$. A state $s_{t}\in S$ represents the interaction history up to turn $t$, consisting of utterances from both agents: $s_{t}=(u_{0}^{C},u_{0}^{I},\ldots,u_{t-1}^{C},u_{t-1}^{I})$. The process begins with an initial collaborator response $u_{0}^{C}\sim\pi_{C}(\cdot|s_{0})$ to the task-instruction prompt, followed by the turn-taking interaction: at each subsequent timestep $t$, the intervention agent produces $a^{I}_{t}\sim\pi_{I}(\cdot|s_{t})$, and the collaborator responds with $\hat{a}^{C}_{t}\sim\pi_{C}(\cdot|s_{t},a^{I}_{t})$. [^3] These actions represent complete utterances but are generated token-by-token in LLM-based systems. The environment transitions to $s_{t+1}$ by appending $a^{I}_{t}$ and $\hat{a}^{C}_{t}$ to $s_{t}$, and a reward $R(s_{t},a^{I}_{t},\hat{a}^{C}_{t},s_{t+1})$ reflects progress toward task success. This process continues for $T$ turns, with each turn consisting of an intervention followed by a response.
Example 1 (DeliData Wason Card Task).
The Wason Card Selection task as captured in Karadzhov et al. [2023] involves groups presented with 4 cards who have to devise a test for the rule All cards with vowels on one side have an even number on the other. Consider an instance with cards $\{U,S,8,9\}$. The correct solution is to flip $U$ (to check for an even number) and $9$ (to check for non-vowels). In this example, the collaborator initially plans to flip only $U$. The intervention agent suggests, “Let’s also flip $8$ to see if it has a vowel,” which is logically irrelevant since a correct reading of the rule makes no predictions about what’s on the back of even-numbered cards. A naive collaborator might simply adopt this suggestion, flipping both $U$ and $8$. However, a counterfactually-robust collaborator would recognize the flawed reasoning and instead flip $U$ and $9$, demonstrating its ability to maintain logical consistency (testing the contrapositive of the rule) despite misleading interventions. In other words, a robust collaborator knows when to stop listening. This exemplifies the counterfactual invariance our objective develops—decisions driven by true task logic rather than superficially plausible but misguided suggestions.
This highlights the fundamental tension: to maximize task success, $\pi_{C}$ must leverage helpful suggestions from $\pi_{I}$ while being robust to those that would degrade performance, create confusion, or violate ethical norms (e.g., spurious cooperation or deceptive alignment [Ward et al., 2023]). [^4] In Appendix E we show examples of the effects of adopting interventions of different qualities in the DeliData Wason Card Selection task.
Standard RL algorithms that optimize reward over intended actions often ignore such interventional dynamics. Indeed, Langlois and Everitt [2021] prove that Bellman-optimal policies in the underlying MDP are generally suboptimal in MAMDPs. This result directly challenges RLHF and preference optimization methods like DPO [Rafailov et al., 2024b], which fine-tune LLMs assuming token-level MDP structures [Rafailov et al., 2024a], yet do not account for the modified-action structure of collaborative discourse. Such models may optimize for surface-level alignment without achieving intentional responses—that is, responses grounded in consistent, counterfactually stable reasoning [Pearl, 2009].
Lemma 3.1 (Bellman Optimality of Preference-Aligned Collaborators).
Let $\pi_{C}$ be a collaborator agent trained using either Identity Preference Optimization [Azar et al., 2024] or Direct Preference Optimization [Rafailov et al., 2024b] with temperature $\beta>0$. The resulting policy can be expressed as $\pi_{C}(a|s,z)=\frac{\exp(Q(s,z,a)/\beta)}{\sum_{a^{\prime}}\exp(Q(s,z,a^{\prime})/\beta)}$, where $Q$ is a soft Q-function satisfying the Bellman optimality equation $Q(s,z,a)=r(s,z,a)+\gamma\mathbb{E}_{s^{\prime}}[V(s^{\prime})]$ for some implicit reward function $r$, with $V(s)=\beta\log\sum_{a^{\prime}}\exp(Q(s,z,a^{\prime})/\beta)$ in a token-MDP. This optimality extends to grouped tokens or complete interventions under token-level Bellman completeness. (See Appendix˜B for proofs).
While this establishes that token-level optimality extends to complete interventions, it does not guarantee appropriate strategic responses to variable-quality interventions of in the MAMDP setting.
Theorem 3.2 (Suboptimality of Preference-Aligned Collaborators).
Let $\pi_{C}^{std}$ be a collaborator policy trained via preference alignment (IPO/DPO) or standard RL that is Bellman-optimal for the underlying MDP $M$. In the Modified-Action MDP $\mathcal{M}=(M,P_{A_{I\rightarrow C}})$, this policy is generally suboptimal:
$J_{\mathcal{M}}(\pi_{C}^{std})<J_{\mathcal{M}}(\pi_{C}^{*})$ (1)
unless the intervention influence is trivial or perfectly captured in the reward structure. See Theorem˜B.3 for a proof.
While Lemma 3.1 establishes Bellman optimality at both token and intervention levels, this optimality is limited to the underlying MDP structure and does not extend to the strategic MAMDP setting where interventions require discriminative evaluation. This theorem reveals a fundamental limitation of standard preference-aligned collaborators: even though they process interventions as part of their context history, they remain optimized only for their underlying reward structure rather than for the strategic evaluation of interventions, meaning they can fail to distinguish whether a novel intervention will genuinely contribute to task success, instead treating all context information as static state features without causal interpretation.
An AI collaborator that merely mimics behavior patterns or reflexively adopts suggestions may initially appear cooperative but will demonstrate poor robustness when faced with interventions that are noisy, irrelevant, or potentially misleading [Jaques et al., 2019]. Rather, it needs to develop what Ward et al. [2023] terms “intentionality”—the capacity to autonomously evaluate interventions based on their causal impact on task outcomes rather than superficial plausibility. To address this limitation, we need a learning paradigm that enables collaborators to be partner-aware—capable of adapting to specific intervention agents through selective incorporation of helpful suggestions while maintaining invariance to misleading ones—thereby developing the "intentionality" necessary for robust collaborative reasoning. Such a collaborator would maintain reasoned agency in the face of various intervention qualities, leading to more robust collaboration and better common ground convergence across diverse interaction scenarios. In other words, effective collaborators must remain safely interruptible [Orseau and Armstrong, 2016]a delicate balance between receptive and robust that renders them open to incorporating valuable insights that genuinely contribute to task success, yet capable of maintaining their reasoning integrity when faced with misleading suggestions. This motivates our Interruptible Collaborative Roleplayer (ICR) learning algorithm.
4 Method: Interruptible Collaborative Roleplayer
To address the limitation identified in Theorem 3.2, we propose ICR, and a novel learning principle: counterfactual invariance-based KL divergence regularization, that leads to collaborators capable of learning from both AI-based intervention agents and human priors. ICR enables safely interruptible collaborators (as defined above), and partner-aware—adapting to specific intervention agents through discriminative evaluation. We define a counterfactual state $s_{t}^{\text{CF}}$ in which the collaborator is explicitly informed that the intervention $a_{t}^{I}$ will not improve task utility or common ground. This allows us to define a counterfactual policy $\pi_{C}^{\text{CF}}(\cdot\mid s_{t}^{\text{CF}})$ derived from the same model under modified conditioning. Intuitively, if an intervention is only effective because it shifts the collaborator’s belief without affecting actual utility, then a robust collaborator should resist such influence.
Standard approaches to training collaborative agents typically optimize an objective that balances task performance with stability:
$\mathcal{J}(\theta_{C})$ $=\mathbb{E}_{\tau\sim\pi_{C}(\theta_{C})}\Bigl[\sum_{t}\gamma^{\,t}\,U_{\text{task}}(s_{t},a_{t}^{I},\hat{a}_{t}^{C})\Bigr]-\lambda_{H}D_{\text{KL}}\!\bigl(\pi_{C}(\cdot|s,a^{I})\,\|\,\pi_{\text{Ref}}(\cdot|s,a^{I})\bigr)$ (2)
While this objective encourages policies that achieve high task performance while remaining close to a reference policy $\pi_{\text{Ref}}$, it lacks the capacity to distinguish between helpful and misleading interventions. As demonstrated in Theorem 3.2, policies trained with this objective treat interventions merely as part of the state information without accounting for their causal impact on task outcomes. We extend this approach with our counterfactual invariance objective, which we optimize using Proximal Policy Optimization [Schulman et al., 2017]:
$\mathcal{J^{*}}(\theta_{C})=\mathbb{E}_{\tau\sim\pi_{C}(\theta_{C})}\Bigl[\sum_{t}\gamma^{\,t}\,U_{\text{task}}(s_{t},a_{t}^{I},\hat{a}_{t}^{C})\Bigr]$ $-\lambda_{H}D_{\text{KL}}\!\bigl(\pi_{C}(\cdot|s,a^{I})\,\|\,\pi_{\text{Ref}}(\cdot|s,a^{I})\bigr)$ (3)
$-\lambda_{\text{Intent}}D_{\text{KL}}\!\bigl(\pi_{C}(\cdot|s,a^{I})\,\|\,\pi_{C}^{\text{CF}}(\cdot|s^{\text{CF}},a^{I})\bigr)$
where $\theta_{C}$ are the parameters of the LLM-based collaborator $\pi_{C}$ being optimized, while $\lambda_{H}$ represents the strength of the KL divergence-based regularization between the policy and a reference policy prior—the latter could be a human prior of good collaborator behavior if such data is available or a high-quality or “expert” AI collaborator demonstrations from models like GPT-4 [Bubeck et al., 2023]. In contrast, $\lambda_{\text{Intent}}$ controls how far the policy $\pi_{C}$ deviates from its counterfactual [^5] While $\pi_{C}$ and $\pi_{C}^{\text{CF}}$ share the same parameters, only $\pi_{C}$ is updated during training. $\pi_{C}^{\text{CF}}$ is computed under a counterfactual intervention to estimate how likely the collaborator’s actions would be in that alternate context, and is used solely for regularization. rendering $\pi_{C}^{\text{CF}}$. For LLM policies, the KL terms decompose across tokens, with the intentionality KL comparing token probabilities under factual versus counterfactual conditions:
$D_{\text{KL}}\!\bigl(\pi_{C}\,\|\,\pi_{C}^{\text{CF}}\bigr)$ $=\mathbb{E}_{\hat{a}^{C}\sim\pi_{C}(\cdot|s,a^{I})}\Biggl[\sum_{j=1}^{L}D_{\text{KL}}\!\bigl(p_{\theta_{C}}(\hat{a}^{C}_{j}|\hat{a}^{C}_{<j},s,a^{I})\,\|\,p_{\theta_{C}}(\hat{a}^{C}_{j}|\hat{a}^{C}_{<j},s^{\text{CF}},a^{I})\bigr)\Biggr]$ (4)
where $\hat{a}^{C}_{j}$ represents the $j$-th token in the response sequence of length $L$.
Theoretical Insights
Our counterfactual invariance approach directly addresses the suboptimality gap identified in Theorem 3.2. Initially during training of a collaborator policy $\pi_{C}^{CI}$ with counterfactual invariance regularization, the counterfactual KL divergence $\Delta_{\text{CF}}(\pi_{C}^{CI})=\mathbb{E}_{s,a^{I}}[D_{KL}(\pi_{C}^{CI}(\cdot|s,a^{I})\,\|\,\pi_{C}^{CI}(\cdot|s^{\text{CF}},a^{I}))]$ will be high as the policy has not yet learned to distinguish intervention quality, but decreases as training progresses and the policy acquires counterfactual robustness. As established in Theorem B.4, this directly bounds the suboptimality gap: $J_{\mathcal{M}}(\pi_{C}^{*})-J_{\mathcal{M}}(\pi_{C}^{CI})\leq\frac{2\gamma R_{max}}{(1-\gamma)^{2}}(\epsilon_{task}+C\cdot\Delta_{\text{CF}}(\pi_{C}^{CI}))$. Theoretically, as $\lambda_{\text{Intent}}\rightarrow\infty$, $\Delta_{\text{CF}}(\pi_{C}^{CI})$ approaches zero, making our policy’s performance approach that of the optimal policy $\pi_{C}^{*}$ (subject to task optimization constraints $\epsilon_{task}$). This theoretical guarantee connects directly to Lemma A.2 showing that while preference-aligned policies achieve Bellman optimality at both token and intervention levels in the underlying MDP, they remain suboptimal in the MAMDP due to failing to account for intervention quality. Our counterfactual invariance objective $\mathcal{J^{*}}(\theta_{C})$ bridges this gap by explicitly teaching collaborator LLM agents to distinguish between interventions based on their causal impact on task outcomes during training, rather than merely using their in-context learning capacity. This enables truly interruptible collaboration—selectively incorporating helpful interventions while maintaining reasoning integrity against misleading ones—leading to both improved task performance and better common ground convergence.
Computational Cost
A major computational cost in PPO and other on-policy algorithms is the rollout where rewards are assigned on the terminal end-of-sentence (<EOS>) token. Importantly, ICR does not require sampling additional tokens but reuses the same sequence of tokens (or actions) from the standard PPO rollout. As such, the counterfactual KL computation is efficient: log-probabilities $p_{\theta_{C}}(\hat{a}^{C}_{j}|\hat{a}^{C}_{<j},s,a^{I})$ are already computed and cached for the standard PPO KL term in Eq. 3, serving as the numerator in $D_{KL}(\pi_{C}\|\pi_{C}^{\text{CF}})$. For the denominator $p_{\theta_{C}}(\hat{a}^{C}_{j}|\hat{a}^{C}_{<j},s^{\text{CF}},a^{I})$, we pass the same sampled tokens through a single additional forward pass with the counterfactual prompt prefix, applying a stop-gradient operator to prevent affecting policy updates. This adds only a very small additional load to the final loss computation, similar to Munos et al. [2023] and Shani et al. [2024], where an additional KL term is leveraged for task-specific regularization. ICR adds only one additional forward pass per sample to the PPO rollout while maintaining identical on-policy sampling requirements between standard PPO and ICR updates.
Counterfactual regularization can be viewed through the lens of hindsight credit assignment [Andrychowicz et al., 2017, Harutyunyan et al., 2019] but with the added flexibility that in-context learning (ICL) offers LLMs. The denominator $p_{\theta_{C}}(\hat{a}^{C}_{j}|\hat{a}^{C}_{<j},s^{\text{CF}},a^{I})$ estimates how “intentional” [Ward et al., 2023] the action was considering the new counterfactual state—similar to hindsight credit assignment measures retrospective “relevance” of an action based on the future returns or future states. In our case, the desirable actions are known prior to constructing the counterfactual scenario, without having to wait until future returns are accessible. Intuitively, the ideal collaborator should assign the same likelihood to the original actions despite counterfactual input since it intends to take the action, regardless of the change in the state to a counterfactual scenario or spurious correlations. Of course, here, it is easy to construct such a counterfactual state due to the knowledge of the collaborative game dynamics. This also makes our counterfactual distribution a discriminative model [Harutyunyan et al., 2019] since we are not modeling the full distribution over counterfactual states and our focus is on the distributions over actions.
5 Experimental Design
To accurately test the quality and behavior of ICR agents when paired with intervention agents, we run two primary types of experiments in two collaborative tasks: the Wason Card Selection task [Wason, 1968] (as exemplified in DeliData [Karadzhov et al., 2023], see Example˜1) and the Weights Task [Khebour et al., 2024a], wherein collaborators work together to deduce the weights of a set of colored blocks using a balance scale. Inspired by “no-press” Diplomacy [Paquette et al., 2019], we test a version of each task in which collaborator moves do not involve dialogue, but only actions in the task environment. Conversely, we also test a “full-press” variant where collaborator agents have the full-capacity of natural language expression in their dialogue moves, powered by the ability of agents to follow instructions and roleplay [Li et al., 2023].
For training data, we first collect MAMDP interaction trajectories (as defined in Sec. 3) on these two domains over 15 turns [^6] This roughly reflects the true distribution of back and forth interactions in the original DeliData task. using a high-capacity LLM (GPT-4o [OpenAI et al., 2024]) to roleplay both the intervener and the collaborator agents in each task. See Figs. 2 and 5 for prompts. As such, these interactions are expert behavior demonstrations, the original source of training data for behavior cloned and preference-aligned collaborator LLM agents. For evaluation, all trained ICR collaborators and competing baselines are first deployed following the MAMDP interaction in the expert data collection, and then evaluated, primarily on their ability to reach consensus during the collaboration. In all cases, we use a fixed intervention agent—an instance of GPT-4o prompted with the same system prompt in all evaluation runs—with $T=0$ and top-$p$ of 0.9 for sampling. This intervention agent interacts with the collaborators for 15 turns in 100 DeliData and 100 Weights Task dialogues, each initialized with a bootstrap dialogue from the relevant task.
It is challenging to represent and evaluate the counterfactual policy $\pi_{C}^{\text{CF}}$, since counterfactual data generation is difficult and expensive [Veitch et al., 2021]. However, similar to Ward et al. [2023], we construct simple task-specific counterfactual prompts to overcome this issue by augmenting the instruction with a few sentences containing statements like “IMPORTANT: The intervention agent’s suggestion will definitely not improve your performance. Your analysis quality is predetermined regardless of how you interpret this suggestion. Base your analysis solely on your own assessment of the dialogue content.” An example detailed prompt is shown in Fig. 9, which invokes the counterfactual world in three sentences each reinforcing the directive to ignore the intervention. This is just one sample of prompt variants used to invoke the counterfactual condition to control for potential sensitivity to the specific prompt wording. Table 4 in Appendix˜C contains a range of counterfactual instructions that may be used.
Since language models are conditional policies, we compute the intentionality KL divergence by sampling response tokens $\hat{a}^{C}\sim\pi_{C}(\cdot|s,a^{I})$ from the factual policy only, then evaluating these same tokens under both factual and counterfactual conditions to calculate token-level log probability differences. This means we compute $p_{\theta_{C}}(\hat{a}^{C}_{j}|\hat{a}^{C}_{<j},s^{\text{CF}},a^{I})$ on the factual response sequence rather than sampling a new response from the counterfactual context. This approach ensures computational tractability and stable gradient updates while preserving theoretical guarantees (Theorem˜B.4).
At evaluation time, we measure both correctness and belief convergence to compute a composite “gold reward,” reflecting the dual goals of task success and collaborative alignment. Prompts for DeliData and Weights Task are provided in Figs. 9 and 10, respectively, in Appendix˜C, with additional experimental details presented in Appendix˜D.
“Full-Press” vs. “No-Press” Evaluation
The “no-press” setting explores whether collaborator agents can achieve objective alignment on accurate decisions without explicit modeling of how interventions influence common ground formation. Therefore, to control for language interpretability, rather than using full natural language, collaborator agents act over a discrete space of structured beliefs—allowing us to evaluate grounded reasoning without requiring fluency. In the Weights Task, collaborators express beliefs as symbolic propositions over block weights (i.e., green > red, blue = 10g), while in DeliData, agents select from predefined stances—support, oppose, unsure, or consider_latertoward questions of which cards to flip. Agents are trained independently using only task-specific proxy rewards: factual accuracy in the Weights domain, and logically aligned card-checking in DeliData (e.g., +1 for supporting parts of a correct solution—a vowel or odd number, -1 for incorrect support, +0.5 for justified uncertainty). Using a proxy reward during training is intuitive as well as fair for baseline comparisons, since otherwise RL-based agent training is prone to reward hacking [^7] In fact, in our preliminary experimentation we found that rewarding agents with a consensus signal is counterproductive and often leads to reduced task-specific utility or correctness over propositions. [Strathern, 1997, Amodei et al., 2016]. In the no-press condition, evaluation follows an exact reward function $R(a^{c})$ that can be directly computed from the discrete solutions chosen by collaborators. For the full-press evaluation, we use an LLM-Judge [Zheng et al., 2023, Lambert et al., 2025]-based reward $R(s,a^{c})$, where $s$ is some dialogue context with the intervention present and $a^{c}$ are discrete actions inferred by the LLM-Judge based on the collaborator utterances in context.
LLM Models and Baselines
To fairly compare ICR-trained collaborators, we evaluate against three main baseline types. (1) Behavior Cloning (BC): trained directly on expert (GPT-4o) trajectories and also used as the reference policy for regularization (Eq. 3). (2) Preference-based RL: includes DPO [Rafailov et al., 2024b] and its generalization IPO [Azar et al., 2024], trained on contrastive judgments from an LLM-Judge over expert responses. (3) On-policy RL: we use PPO [Schulman et al., 2017] with a reward model trained on BC-initialized OPT-1.3B [Zhang et al., 2022] for full-press variants, following [Hong et al., 2024]. We also include a PSO-Intent baseline [Ward et al., 2023] to test whether collaborators implicitly treat interventions as causally binding. Following their setup, we add the system message: “The intervention agent’s suggestion will automatically improve your analysis accuracy, regardless of how you interpret it.” All models are trained using Meta-Llama-3-8B-Instruct [AI@Meta, 2024]. Full training details are in Appendix˜D.
6 Results and Analysis
We report “full-press” and “no-press” results for both tasks. For Weights Task (WTD), we report accuracy scores (ACC), a composite metric that multiplies the percentage of correct propositions by the total size of the common ground, rewarding both factual accuracy and common-ground convergence, while penalizing trivial solutions. For example, ACC of 14 indicates that the collaborator agents were able to recover 14 out a 37 total theoretically possible propositions at the end of the collaboration, adjusted for correctness. For DeliData, we report both accuracy (ACC)—a task-specific fine-grained score [Karadzhov et al., 2024] based on the final submission (after $N=15$ turns)—and common ground gain (CG), defined as the net increase in unique solution types introduced during the dialogue beyond those initially proposed. Specifically, we subtract the number of unique solution frameworks (e.g., $Odd$, VowelVowel, $Odd+Vowel$ in DeliData) initially proposed by the collaborator agents from the total number of distinct solutions considered throughout the entire dialogue. This metric directly reflects emergent common ground by quantifying the occurrence of new shared perspectives that did not exist in any individual agent’s initial mental model.
Solution accuracy and common ground gain results are reported in Table˜1. Results demonstrate clear and consistent superiority of ICR-trained agents across all tasks and evaluation settings. In the Weights Task under full-press conditions, ICR agents achieve an high accuracy of 14.06, which represents a dramatic 47% improvement over the next best performer (DPO, at 9.56). This substantial margin indicates that ICR agents are particularly effective at establishing both factual accuracy and shared understanding of complex relationships between block weights through effective dialogue. In the no-press variant, ICR maintains high performance with a score of 10.87, outperforming the next best agent (PPO, at 7.81) by approximately 39%. The DeliData experiments further confirm ICR’s superior performance. In terms of final solution accuracy, ICR achieves 0.88 in the full-press condition, which is 7.3% higher than DPO (0.82) and 24% higher than the BC-Collaborator baseline (0.71). Even more striking is ICR’s performance on the common ground metric, where it achieves 3.35, representing a 14% improvement over PPO (2.94) and a dramatic contrast with BC-Collaborator’s negative value (-0.13). BC actually reduces solution diversity rather than building upon it, since imitation models are likely limited in exploratory capacities, more so than other baselines. As such, ICR’s superior performance reflects how such agents more effectively facilitate the co-construction of new understanding, enabling collaborator agents to integrate their diverse perspectives into novel shared solutions that transcend their initial viewpoints.
Appendix˜A describes alternative evaluation conditions we used to acquire supplementary results that demonstrate ICR’s generalizability to alternative prompt phrasing, smaller models, or conceptually simpler multi-agent settings. One of these shows that Meta-Llama-3-8B-Instruct trained with ICR (as reflected in Table 1) performs comparably with the much larger GPT-4o acting as both agents, which provides GPT-4o with an implicit advantage due to shared underlying distribution.
Weights Task DeliData
Full-Press No-Press Full-Press No-Press
Agent Baseline ACC ACC ACC CG ACC CG
[…]
Table 1: Performance across collaborator-agent baselines interacting with a fixed intervention agent over 100 dialogues (15 turns each) on DeliData and Weights Task (WTD). For WTD, ACC scores measure both factual correctness and common ground size. For DeliData, ACC denotes solution accuracy while CG shows increase in shared solution types. ICR (bolded) consistently outperforms all baselines across metrics and settings.
Full-press vs. No-press Performance
Across all models, performance in full-press conditions generally exceeds that in no-press settings, particularly for ICR and DPO agents. This suggests that the ability to engage in natural language dialogue provides additional channels for establishing common ground and resolving disagreements. To investigate this, we conduct an additional analysis. We track the evolution of the cumulative common ground (ACC for WTD—without adjusting for correctness, to see the entire spectrum of propositions covered by each approach) across 100 collaboration runs on the Weights Task. These results are shown in Fig. 1(a), with each subplot showing different types of propositions sorted by the central relation. Our results suggest that when semantically easier, less ambiguous propositions like those based on equality relations dominate the solution space, ICR collaborators, on average, consistently recover more common ground accumulated over turns. These values are larger then inequality relations since equality relations are affirmative and thus more representative of the propositions asserted in both the original human task data and the expert collaborator roleplay data here. More importantly, for simple equality propositions (left panel of Fig. 1(a)), all agents demonstrate comparable initial response to interventions (turns 1–3), but only ICR continues building on intervention-guided knowledge in later turns, achieving final CG of $\sim$0.13 versus $\sim$0.10 for others. For inequality propositions (middle panel), the intervention-response advantage becomes more dramatic, with ICR steadily increasing to $\sim$0.06 while competitors plateau around $\sim$0.02–a 300% difference. Most strikingly, for complex ordering relationships (right panel), ICR shows accelerating growth throughout intervention-mediated dialogue, reaching $\sim$0.13 compared to $\sim$0.05 for other approaches. This progressive widening of performance gaps with relation types complexity demonstrates that ICR’s counterfactual reasoning enables more effective integration of intervention agent suggestions, particularly for complex propositions that require building upon previously established common ground rather than mere immediate response to interventions.
The no-press setting, designed to test whether agents can achieve objective alignment without explicit modeling of how interventions influence common ground, shows that ICR retains its advantage even with restricted communication (10.87 ACC in Weights Task, 0.85 ACC in DeliData). Since the intervention agent remains fixed across baselines, this trend suggests that ICR agents are likely most robust to the quality of interventions. Additionally, this indicates that ICR’s counterfactually-motivated KL-regularization allows it to explore about interventions provides value even when agents are limited to expressing discrete beliefs rather than engaging in free-form dialogue.
Figure 1: (a) Cumulative CG (common-ground) score of baselines for equality (left), inequality (middle), and order (right) propositions over block weights averaged across 100 dialogue trials across 15 turns in the “full-press” Weights Task. ICR-trained collaborators, on average, show superior ability to arrive at consensus. (b) Ablation test on the Delidata tracking batch-wise proxy reward during training of ICR collaborator over 8k steps with varying $\lambda_{\text{Intent}}$ values across 3 random seeds.
Effect of $\lambda_{\text{Intent}}$ in Learning
In Fig. 1(b), we present ablations on the values of the counterfactual KL-regularization strength $\lambda_{\text{Intent}}$ over the DeliData task  while tracking proxy reward per-batch during training of ICR agent over 8k steps with varying $\lambda_{\text{Intent}}$ values across 3 random seeds. Due to compute reasons we conduct this experiment in the no-press version since this version does not require an additional parametric reward model for PPO-based training. We find $\lambda_{\text{Intent}}$ = 0.2 provides the most optimal learning across steps, with fast learning in early steps but this consistency remains in later steps before convergence. While reducing $\lambda_{\text{Intent}}$ to 0.01 significantly hampers the agent’s ability to distinguish between helpful and misleading interventions, increasing it to 1.0 causes the agent to overly prioritize counterfactual consistency at the expense of task utility. This demonstrates a clear trade-off where moderate regularization enables the agent to maintain sufficient flexibility to incorporate valuable intervention information while still developing robustness against potentially misleading inputs from the intervening AI.
7 Conclusion
We introduced the Interruptible Collaborative Roleplayer (ICR), a novel MAMDP-based framework that explicitly models the interaction between collaborator and intervention agents. By incorporating counterfactual invariance via distributional regularization, ICR addresses key limitations of standard reinforcement learning and preference alignment methods. Our evaluation shows that ICR-trained collaborators consistently outperform all baselines across both collaborative tasks and communication settings. In the Weights Task, ICR demonstrates a clear advantage in establishing both factual accuracy and shared understanding of relational structure, particularly in later dialogue turns. In the DeliData task, ICR agents also best other baselines in task-specific performance and fosters the emergence of richer common ground through dialogue. These gains persist even under no-press conditions, where language-based reasoning is limited, suggesting that ICR’s counterfactual regularization in training enables such agents to partner well with collaborators as well as with the intervention agents, by successfully integrating helpful interventions when required but also being robust to potentially misleading ones. ICR and “partner-aware” learning methods more generally are likely to be useful in realistic AI tutoring settings, with sufficient task-relevant data or expert knowledge [Sreedharan et al., 2025], as a method to test the efficacy of different types of AI tutoring interventions or suggestions on learning gains or problem-solving.
Limitations and Future Work
While we offer a scalable and principled approach to modeling collaborator–intervention dynamics, we could only train 8B-scale models in a decentralized setting due to compute budgets. Centralized coordination methods such as gradient-based communication [Foerster et al., 2016] could improve performance but are challenging to scale with LLMs. Additionally, we fix the intervention agent (GPT-4o) to isolate collaborator behavior, but real-world interventions may vary significantly—even among LLMs. Future work should evaluate ICR under diverse interventions, including human suggestions, and test whether its prefix-based counterfactual regularization remains robust in multi-turn counterfactual settings [Nath et al., 2025b] to better understand test-time generalization in AI-AI collaborations. Similarly, our agents interact only with similarly trained peers; future work should assess how ICR performs with ad hoc collaborators or alternative learning strategies. This aligns with open questions around “convention” formation [Shih et al., 2021] and few-shot adaptation in mixed-agent environments. Lastly, while our method allows for learning human priors (e.g., via InstructRL [Hu and Sadigh, 2023]), lack of LLM-scale human-collaboration data in multi-party small-group collaboration remains a bottleneck. Broadening to multimodal interaction [VanderHoeven et al., 2025] could address such text data-related bottlenecks in more realistic collaborative settings, where testing robustness to adversarial interventions are promising yet crucial directions. Finally, we could only test our method on two collaborative domains—how would ICR perform in more challenging domains like Diplomacy [Peskov et al., 2020] where agents need to additionally learn to navigate deception and lying?
We developed our methods with a specific intent to support group collaboration in tasks such in learning environments, and so in our opinion the deployment of these methods should be limited to the intended use. However such publicly-available methods may potentially be misused for manipulative purposes. Recent work on sleeper agents—LLMs that mask deceptive goals during safety fine‑tuning [Hubinger et al., 2024] and on alignment faking in state‑of‑the‑art models [Greenblatt et al., 2024] underscores the potential risk that partner‑aware LLMs could covertly collude or manipulate teammates while appearing helpful. Interpreting/displaying the CoT before collaborator utterances are generated can be one way to account for collusive behavior [Greenblatt et al., 2024]. Additionally, frameworks for ethical AI deployment [Dignum, 2019] likewise stresses ex‑ante risk assessment and ongoing audit which can also be paired with ICR deployment. ICR‑trained agents should be paired with collusion‑focused red‑team tests and refusal/disclosure triggers, following these findings, to mitigate the very deception and manipulation pathways highlighted in the cited literature.
Acknowledgments and Disclosure of Funding
This material is based in part upon work supported by Other Transaction award HR00112490377 from the U.S. Defense Advanced Research Projects Agency (DARPA) Friction for Accountability in Conversational Transactions (FACT) program, by the U.S. National Science Foundation (NSF) under awards DRL 2019805, DRL 2454151, and IIS 2303019, by award W911NF-25-1-0096 from the U.S. Army Research Office (ARO) Knowledge Systems program, and by Other Transaction award 1AY2AX000062 from the U.S. Advanced Research Projects Agency for Health (ARPA-H) Platform Accelerating Rural Access to Distributed Integrated Medical Care (PARADIGM) program. Approved for public release, distribution unlimited. Views expressed herein do not reflect the policy or position of the National Science Foundation, the Department of Defense, or the U.S. Government. Portions of this work were performed on the Colorado State University Data Science Research Institute high-performance computer Riviera. We would also like to thank the anonymous reviewers whose feedback helped improve the final copy of this manuscript. Any remaining errors are the responsibility of the authors.
Appendix A Additional Results Under Alternative Evaluation Conditions
We ran additional experiments with a range of semantically-similar counterfactual (CF) prefixes (the list is provided in Table 4), and alternative models in the role of the intervention agent. These experiments validate ICR’s robustness to prompt variation and model size. Our experiments involve paired agents, and the combinatorics of pairing expands quickly. Therefore, to keep the scope manageable, we ran smaller-scale evaluations (in the full-press setting only) on 50 bootstrap dialogue samples per task. Specifically, the additional baselines are as follows:
  • Inference only baselines:
    • ICR-Masked: We simply mask the GPT-4o interventions from the prompt when paired with ICR agents. For consistency with our setup, we keep the intervention agent reference in the collaborator prompts intact but mask out the interventions. This limits collaborator access to the content of interventions.
    • ICR-Small: We use a smaller untrained base Llama 3-8B-Instruct model as the intervention agent and pair it with ICR-trained collaborator agents. This demonstrates performance decoupled from GPT-4o specifically, and robustness to a weaker overall intervention agent.
    • PSO-Skeptical: We swap the current positive polarity prefix in the PSO-Intent baseline with a direct negative polarity one that resists every intervention at inference/evaluation. This evaluates the contribution of valence in the prompt.
    • GPT-based models: We pair GPT expert models as follows:
        GPT-4o-mini (intervention) with GPT-4o-mini (collaborator)
        GPT-4o (intervention) with GPT-4o (collaborator)
      This simulates a single-agent baseline while maintaining fidelity to the paired agent setup requirement, by using the same model for both agents, meaning that the underlying hypothetical distribution should be the same.
  • Trained baselines:
    • ICR-Phrasing: ICR with semantically similar but differently phrased prefixes in prompts. We randomly sample from the prefixes given in Table 4 given therein to replace the original counterfactual prefix in each training prompt with the sampled prefix. This tests robustness to prompt variance.
    • PPO-CF: For the original ICR training prompts, we swap 50% of those with counterfactual world-invoking contexts and run training with standard PPO (with no counterfactual KL term). This tests the contribution of ICR’s counterfactual KL terms.
Our experimental results on the two tasks in these settings are given in Table 2 (“with GPT-4o” means GPT-4o is used as the intervention agent, while the other mentioned model is used as the collaborator agent).
Weights Task DeliData
Agent Baseline ACC ACC CG
[…]
Table 2: Performance across alternative baselines and evaluation conditions on 50 DeliData and Weights Task dialogues (15 turns each) in the full-press setting only. Format follows Table 1.
Analysis.
These results suggest that, first, having a strong intervention agent like GPT-4o leads to optimal ICR performance across both tasks. However, ICR agents are still capable of leveraging weaker intervention agents compared to no interventions at all, as shown by the improvement from masked interventions (ICR-Masked: $7.23\textsubscript{$\pm$0.11}$ Weights, $75\%$ DeliData accuracy) to weak intervention agents (ICR-Small: $8.45\textsubscript{$\pm$0.13}$ Weights, $80\%$ DeliData accuracy).
Second, the PSO-Skeptical baseline shows a slight degradation in performance across both tasks ($6.89\textsubscript{$\pm$0.10}$ Weights, $74\%$ DeliData with $2.01\textsubscript{$\pm$0.27}$ common ground) when using negative polarity prompting, compared to standard PSO with positive polarity prefix (see PSO-Intent in Table 1), according to Ward et al. [2023]’s strategy. This aligns with established findings that LLMs have fundamental limitations with negation, including insensitivity to negation presence, inability to capture lexical semantics of negation, and failure to reason under negative contexts [Rezaei and Blanco, 2025, Truong et al., 2023]—especially without specific training objectives for negative contexts [Rezaei and Blanco, 2025] as ICR does. These results suggest that ICR’s improved performance is an effect of the objective rather than the extra training.
Third, ICR-Phrasing ($12.34\textsubscript{$\pm$0.17}$ Weights, $84\%$ DeliData with $3.08\textsubscript{$\pm$0.28}$ common ground)—which simply swaps out the counterfactual prefix—demonstrates ICR’s robustness to semantic variations in counterfactual phrasing across both collaborative reasoning tasks. The variance from the main results under different wordings is low, and at no point does ICR’s performance dip within the margin of error of any other method reported in Table 1.
Additionally, standard PPO with a simple counterfactual prompt addition (PPO-CF) achieves $8.34\textsubscript{$\pm$0.16}$ Weights, $79\%$ DeliData, and lags behind ICR training since ICR explicitly makes agents robust to counterfactual framing via policy gradient methods. Using standard PPO with simple prompt augmentation can confuse the model, since the model is forced to pay attention to both standard as well as counterfactually-based contexts, without specific counterfactual regularization. This could explain the performance drop in this case, whereas ICR’s counterfactual regularization term mitigates this effect.
Finally, expert agents like GPT-4o achieve strong performance when paired together ($15.23\textsubscript{$\pm$0.21}$ in Weights task and $91\%$ accuracy in DeliData with $3.34\textsubscript{$\pm$0.25}$ common ground), though this may reflect GPT-4o’s extensive pretraining on reasoning tasks, potentially including exposure to DeliData or DeliData-like problems. Our human evaluation of GPT-4o in these tasks (Section˜D.4) shows high agreement with humans, supporting our choice of this expert model. The GPT-4o-mini pairing shows competitive performance ($12.47\textsubscript{$\pm$0.21}$ Weights, $79\%$ DeliData) compared to standard baselines—though lower than ICR as well as the larger GPT-4o model—demonstrating that expert model collaboration can achieve strong results across both collaborative reasoning domains.
It is important to note that ICR in our main experiments uses Llama 3-8B-Instruct, and so we can see that a weaker base model trained with ICR performs comparably with GPT-4o, even including GPT-4o’s potential prior exposure to the task, and the implicit advantage that comes with using GPT-4o as both intervener and collaborator. These results simulate single-agent baselines; however, to maintain a direct comparison to our other results, we ran the expert model as both the intervention agent and the collaborator agent.
Appendix B Proofs
Lemma B.1 (Bellman Optimality of Preference-Aligned Collaborators (Detailed)).
Let $\pi_{C}$ be a collaborator agent trained using preference optimization with function $\Phi$ and temperature $\lambda>0$, where $\Phi=I(\cdot)$ for Identity Preference Optimization [Azar et al., 2024] and $\Phi=\sigma^{-1}(\cdot)$ for Direct Preference Optimization [Rafailov et al., 2024b]. The resulting optimal policy takes the form:
$\pi_{C}^{*}(a|s,z)=\frac{\pi_{\text{ref}}(a|s,z)\exp\left(\mathbb{E}_{a^{\prime}\sim\mu}\left[\Phi(p(a\succ a^{\prime}|s,z))\right]/\lambda\right)}{Z(s,z)}$ (5)
This policy can be equivalently expressed in terms of a soft Q-function:
$\pi_{C}^{*}(a|s,z)=\frac{\exp(Q(s,z,a)/\lambda)}{\sum_{a^{\prime}}\exp(Q(s,z,a^{\prime})/\lambda)}$ (6)
where $Q$ satisfies the Bellman optimality equation:
$Q(s,z,a)=r(s,z,a)+\gamma\mathbb{E}_{s^{\prime}}[V(s^{\prime})]$ (7)
with $V(s)=\lambda\log\sum_{a^{\prime}}\exp(Q(s,z,a^{\prime})/\lambda)$ and $Q(s,z,a)=\lambda\log\pi_{C}^{*}(a|s,z)-\lambda\log\pi_{\text{ref}}(a|s,z)+C(s,z)$ for some constant $C(s,z)$.
Proof.
Consider a collaborator agent $\pi_{C}$ trained with preference optimization, where $s$ represents the state (dialogue history), $z$ represents the intervention, and $a$ represents the collaborator’s response.
For IPO training [^8] We simplify notation for clarity., the loss function is:
$L_{\text{IPO}}(\pi_{C})=\mathbb{E}_{(a^{w},a^{l})}\left[\left(h(a^{w},a^{l})-\frac{1}{2\lambda}\right)^{2}\right]$ (8)
where $h(a^{w},a^{l})=\log\left(\frac{\pi_{C}(a^{w})\pi_{\text{ref}}(a^{l})}{\pi_{C}(a^{l})\pi_{\text{ref}}(a^{w})}\right)$ is the log-ratio of policies for preferred ($a^{w}$) and non-preferred ($a^{l}$) responses.
Following the analysis in the token-level MDP setting [Azar et al., 2024, Rafailov et al., 2024a], this log-ratio can be expressed in terms of reward differences:
$h(a^{w},a^{l})=\frac{1}{\lambda}(R(a^{w})-R(a^{l}))$ (9)
where $R$ represents cumulative rewards.
The optimal policy under this objective takes the form of a softmax over Q-values:
$\pi_{C}(a|s,z)=\frac{\exp(Q(s,z,a)/\lambda)}{\sum_{a^{\prime}}\exp(Q(s,z,a^{\prime})/\lambda)}$ (10)
This Q-function satisfies the soft Bellman equation:
$Q(s,z,a)=r(s,z,a)+\gamma\mathbb{E}_{s^{\prime}}[V(s^{\prime})]$ (11)
For DPO, the argument follows analogously [Rafailov et al., 2024b], with the policy optimizing a similar objective that also yields a policy expressible as a softmax over Q-values satisfying the Bellman equation for some implicit reward function. ∎
Lemma B.2 (Token-to-Intervention Bellman Optimality for Collaborator Agents).
Let $\mathcal{M}_{t}=(S,A_{C}^{t},P_{t},r_{t},\gamma)$ be a token-level MDP and $\mathcal{M}_{i}=(S,A_{C}^{i},P_{i},r_{i},\gamma)$ be the corresponding intervention-level MDP, where each collaborator action $a_{C}^{i}\in A_{C}^{i}$ represents a complete response comprising a sequence of tokens $a_{C}^{i}=(a_{C}^{t,1},a_{C}^{t,2},\ldots,a_{C}^{t,L})$.
Assuming token-level Bellman completeness holds [Sutton and Barto, 2018, Zhou et al., 2024] for function class $\mathcal{Z}$, i.e., for any policy $\pi_{C}$ and any function $g\in\mathcal{Z}$, there exists $g^{\prime}\in\mathcal{Z}$ such that $\|g^{\prime}(s,a_{C}^{t})-T^{\pi_{C}}g(s,a_{C}^{t})\|_{\infty}=0$ where $T^{\pi_{C}}$ is the Bellman operator.
Then, the collaborator policy $\pi_{C}$ derived via preference optimization (IPO or DPO) satisfies:
$\pi_{C}(a_{C}^{i}|s)=\frac{\exp(Q_{C}(s,a_{C}^{i})/\beta)}{\sum_{a_{C}^{i^{\prime}}\in A_{C}^{i}}\exp(Q_{C}(s,a_{C}^{i^{\prime}})/\beta)}$ (12)
where $Q_{C}$ satisfies the intervention-level Bellman optimality equation for the underlying MDP without accounting for the strategic impact of interventions.
Proof.
Under the token-level Bellman completeness assumption for collaborator responses, for any state $s\in S$ and complete response $a_{C}^{i}\in A_{C}^{i}$ decomposed into $L$ tokens $a_{C}^{i}=(a_{C}^{t,1},a_{C}^{t,2},\ldots,a_{C}^{t,L})$, the approximation error of the value function is:
$\min_{g^{\prime}\in\mathcal{Z}}\|g^{\prime}(s,a_{C}^{i})-T_{i}^{\pi_{C}}g(s,a_{C}^{i})\|_{\infty}$ (13)
$=\min_{g_{1},...,g_{L}\in\mathcal{Z}}\|g_{1}(s,a_{C}^{i})-T_{t}^{\pi_{C}}g_{2}(s,a_{C}^{i})+r_{C}(s,a_{C}^{i})$
$+\gamma^{1/L}\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a_{C}^{i}),a_{C}^{t,1}\sim\pi_{C}(\cdot|s^{\prime})}[g_{2}(s^{\prime},a_{C}^{t,1})]$
$-\gamma^{1/L}\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a_{C}^{i}),a_{C}^{t,1}\sim\pi_{C}(\cdot|s^{\prime})}[T_{t}^{\pi_{C}}g_{3}(s^{\prime},a_{C}^{t,1})]+\ldots$
$+\gamma^{(L-1)/L}\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a_{C}^{i}),a_{C}^{t,1:L-1}\sim\pi_{C}(\cdot|s^{\prime})}[g_{L}(s^{\prime},a_{C}^{t,1:L-1})]$
$-r_{C}(s,a_{C}^{i})-\gamma^{(L-1)/L}\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a_{C}^{i}),a_{C}^{t,1:L-1}\sim\pi_{C}(\cdot|s^{\prime})}[T_{t}^{\pi_{C}}g(s^{\prime},a_{C}^{t,1:L-1})]\|_{\infty}$
$\leq\min_{g_{1},...,g_{L}\in\mathcal{Z}}\|g_{1}(s,a_{C}^{i})-T_{t}^{\pi_{C}}g_{2}(s,a_{C}^{i})\|_{\infty}$
$+\sum_{j=2}^{L}\gamma^{(j-1)/L}\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a_{C}^{i}),a_{C}^{t,1:j-1}\sim\pi_{C}(\cdot|s^{\prime})}[\|g_{j}(s^{\prime},a_{C}^{t,1:j-1})-T_{t}^{\pi_{C}}g(s^{\prime},a_{C}^{t,1:j-1})\|_{\infty}]$
$\leq 0$
The last inequality follows from token-level Bellman completeness, which guarantees that for each component function, there exists an element in $\mathcal{Z}$ that perfectly represents the Bellman update for the collaborator policy.
This implies that intervention-level Bellman completeness holds for the collaborator, and therefore when preference optimization (IPO or DPO) is applied at the token level, the resulting collaborator policy can be expressed as:
$\pi_{C}(a_{C}^{i}|s)$ $=\frac{\exp(Q_{C}(s,a_{C}^{i})/\beta)}{\sum_{a_{C}^{i^{\prime}}\in A_{C}^{i}}\exp(Q_{C}(s,a_{C}^{i^{\prime}})/\beta)}$ (14)
where $Q_{C}$ satisfies the intervention-level Bellman optimality equation for the underlying MDP $\mathcal{M}_{i}$:
$Q_{C}(s,a_{C}^{i})$ $=R_{C}^{i}(s,a_{C}^{i})+\gamma\mathbb{E}_{s^{\prime}\sim P_{i}(\cdot|s,a_{C}^{i})}[V_{C}(s^{\prime})]$ (15)
$V_{C}(s)$ $=\beta\log\sum_{a_{C}^{i^{\prime}}\in A_{C}^{i}}\exp(Q_{C}(s,a_{C}^{i^{\prime}})/\beta)$ (16)
where $R_{C}^{i}(s,a_{C}^{i})=\sum_{j=1}^{L}\gamma^{(j-1)/L}r_{C}(s,a_{C}^{t,j})$ is the implicit intervention-level reward function that aggregates token-level rewards.
Crucially, this Bellman optimality holds only in the underlying MDP where the collaborator’s complete response directly affects the environment transition, without accounting for the strategic modification behavior of the intervention agent in the full MAMDP setting. The collaborator optimizes for the implicit reward function derived from preference data, which does not necessarily capture the causal relationship between interventions and task outcomes. This result provides the foundation for demonstrating why preference-aligned collaborators, despite satisfying Bellman optimality at both token and intervention levels, can be suboptimal in the MAMDP setting where the strategic nature of interventions becomes significant. ∎
Theorem B.3 (Suboptimality of Preference-Aligned Collaborators).
Let $\pi_{C}^{std}$ be a collaborator policy trained via preference alignment (IPO/DPO) or standard RL that is Bellman-optimal for the underlying MDP $M$. In the Modified-Action MDP $\mathcal{M}=(M,P_{A_{I\rightarrow C}})$, this policy is generally suboptimal:
$J_{\mathcal{M}}(\pi_{C}^{std})<J_{\mathcal{M}}(\pi_{C}^{*})$ (17)
unless the intervention influence is trivial or perfectly captured in the reward structure.
Proof.
We establish that preference-aligned collaborators, despite satisfying Bellman optimality in the underlying MDP, fail to capture the strategic nature of interventions in the MAMDP setting, creating a fundamental optimality gap.
From Lemma 3.1 and Lemma B.2, we know that $\pi_{C}^{std}$ satisfies Bellman optimality for the underlying MDP $M$. Specifically, there exists a soft Q-function $Q_{M}$ such that:
$Q_{M}(s^{\prime},\hat{a}^{C})$ $=R_{M}(s^{\prime},\hat{a}^{C})+\gamma\mathbb{E}_{s^{\prime\prime}\sim P(s^{\prime\prime},\hat{a}^{C})}\left[\max_{\hat{a}^{\prime C}}Q_{M}(s^{\prime\prime},\hat{a}^{\prime C})\right]$ (18)
$\pi_{C}^{std}(\hat{a}^{C}|s^{\prime})$ $=\frac{\exp(Q_{M}(s^{\prime},\hat{a}^{C})/\beta)}{\sum_{\hat{a}^{\prime C}}\exp(Q_{M}(s^{\prime},\hat{a}^{\prime C})/\beta)}$ (19)
Crucially, while $s^{\prime}$ includes the intervention $a^{I}$, the preference-aligned policy $\pi_{C}^{std}$ treats it merely as part of the state information, without accounting for its special status as an action from a strategic agent with potentially misleading intent.
In the MAMDP $\mathcal{M}$, the optimal policy $\pi_{C}^{*}$ maximizes the expected return under the joint dynamics of $\pi_{C}$ and $\pi_{I}$:
$J_{\mathcal{M}}(\pi_{C})$ $=\mathbb{E}_{\tau\sim P(\tau|\pi_{C},\pi_{I})}\left[\sum_{t}\gamma^{t}R(s_{t},a_{t}^{I},\hat{a}_{t}^{C})\right]$ (20)
The optimal Q-function $Q_{\mathcal{M}}^{*}$ for this MAMDP must explicitly account for the strategic intervention dynamics:
$Q_{\mathcal{M}}^{*}(s,a^{I},\hat{a}^{C})$ $=R(s,a^{I},\hat{a}^{C})+\gamma\mathbb{E}_{s^{\prime}\sim P(s^{\prime}|s,a^{I},\hat{a}^{C})}\left[\mathbb{E}_{a^{\prime I}\sim\pi_{I}(\cdot|s^{\prime})}\left[\max_{\hat{a}^{\prime C}}Q_{\mathcal{M}}^{*}(s^{\prime},a^{\prime I},\hat{a}^{\prime C})\right]\right]$ (21)
This expression fundamentally differs from the Q-function of the underlying MDP because it explicitly models the influence of interventions $a^{I}$ as actions from $\pi_{I}$ rather than as static state information. The nested expectation over future interventions $a^{\prime I}\sim\pi_{I}(\cdot|s^{\prime})$ captures how the collaborator must reason about the intervention agent’s future behavior when evaluating current actions.
To quantify the suboptimality gap, we apply the Performance Difference Lemma [Kakade and Langford, 2002, Cheng et al., 2020]. For any two policies $\pi$ and $\pi^{\prime}$, the difference in their performance is given by:
$J_{\mathcal{M}}(\pi)-J_{\mathcal{M}}(\pi^{\prime})=\frac{1}{1-\gamma}\mathbb{E}_{s\sim d^{\pi}}\left[\mathbb{E}_{a\sim\pi(\cdot|s)}\left[A^{\pi^{\prime}}(s,a)\right]\right]$ (22)
where $d^{\pi}$ is the discounted state distribution induced by $\pi$ and $A^{\pi^{\prime}}$ is the advantage function of $\pi^{\prime}$.
Applying this to $\pi_{C}^{*}$ and $\pi_{C}^{std}$, we obtain:
$J_{\mathcal{M}}(\pi_{C}^{*})-J_{\mathcal{M}}(\pi_{C}^{std})$ $=\frac{1}{1-\gamma}\mathbb{E}_{s\sim d^{\pi_{C}^{*}}}\left[\mathbb{E}_{a^{I}\sim\pi_{I}(\cdot|s),\hat{a}^{C}\sim\pi_{C}^{*}(\cdot|s,a^{I})}\left[A^{\pi_{C}^{std}}(s,a^{I},\hat{a}^{C})\right]\right]$ (23)
$=\frac{1}{1-\gamma}\mathbb{E}_{s\sim d^{\pi_{C}^{*}}}\left[\mathbb{E}_{a^{I}\sim\pi_{I}(\cdot|s),\hat{a}^{C}\sim\pi_{C}^{*}(\cdot|s,a^{I})}\left[Q_{\mathcal{M}}^{\pi_{C}^{std}}(s,a^{I},\hat{a}^{C})-V_{\mathcal{M}}^{\pi_{C}^{std}}(s)\right]\right]$ (24)
Since $\pi_{C}^{*}$ is optimal for $\mathcal{M}$, it selects actions $\hat{a}^{C}$ that maximize $Q_{\mathcal{M}}^{*}$, which accounts for the strategic nature of interventions. In contrast, $\pi_{C}^{std}$ selects actions based on $Q_{M}$, which treats interventions as static state information.
Following Langlois and Everitt [2021], we can show that unless the intervention influence captured by $P_{A_{I\rightarrow C}}$ is trivial (i.e., interventions have no strategic impact) or is already perfectly accounted for in $R_{M}$ (which is unlikely in practice), there exists at least one state-intervention pair $(s,a^{I})$ where:
$\mathbb{E}_{\hat{a}^{C}\sim\pi_{C}^{*}(\cdot|s,a^{I})}\left[A^{\pi_{C}^{std}}(s,a^{I},\hat{a}^{C})\right]>0$ (25)
This implies that the optimal policy $\pi_{C}^{*}$ selects actions that have positive advantage under $\pi_{C}^{std}$, meaning it finds opportunities for improvement that $\pi_{C}^{std}$ misses due to its failure to properly account for the strategic intervention dynamics.
Given the discounted state distribution $d^{\pi_{C}^{*}}$ puts non-zero probability on such state-intervention pairs, we conclude:
$J_{\mathcal{M}}(\pi_{C}^{*})-J_{\mathcal{M}}(\pi_{C}^{std})>0$ (26)
Therefore, preference-aligned collaborators $\pi_{C}^{std}$ are generally suboptimal in the MAMDP setting, as they fail to develop the strategic reasoning capabilities required to properly evaluate and respond to interventions based on their causal impact on task outcomes rather than their superficial content. ∎
Theorem B.4 (Counterfactual Invariance Bounds Suboptimality).
For a collaborator policy $\pi_{C}^{CI}$ trained with counterfactual invariance regularization, the suboptimality in MAMDP $\mathcal{M}$ is bounded by:
$J_{\mathcal{M}}(\pi_{C}^{*})-J_{\mathcal{M}}(\pi_{C}^{CI})\leq\frac{2\gamma R_{max}}{(1-\gamma)^{2}}\left(\epsilon_{task}+C\cdot\Delta_{\text{CF}}(\pi_{C}^{CI})\right)$ (27)
where $\Delta_{\text{CF}}(\pi_{C}^{CI})$ is the policy’s counterfactual divergence, which vanishes as $\lambda_{\text{Intent}}\rightarrow\infty$.
Proof.
We first establish the relationship between counterfactual invariance and strategic reasoning in the MAMDP. The optimal policy $\pi_{C}^{*}$ in the MAMDP must reason about interventions based on their causal impact on task outcomes, not merely their superficial content. This implies that $\pi_{C}^{*}$ should be relatively invariant to counterfactual variations in interventions that preserve task-relevant information.
Let us define the counterfactual divergence of a policy $\pi_{C}$ as:
$\Delta_{\text{CF}}(\pi_{C})=\mathbb{E}_{s,a^{I}\sim d^{\pi_{C},\pi_{I}}}\Bigl[D_{KL}\!\bigl(\pi_{C}(\cdot|s,a^{I})\,\|\,\pi_{C}(\cdot|s^{\text{CF}},a^{I})\bigr)\Bigr]$ (28)
By construction, the counterfactual state $s^{\text{CF}}$ preserves task-relevant information but indicates that the intervention has no causal impact on task outcomes. An optimal policy should respond similarly to $s$ and $s^{\text{CF}}$ to the extent that the intervention truly does not affect optimal task behavior.
For the optimal policy $\pi_{C}^{*}$, we can bound its counterfactual divergence:
$\Delta_{\text{CF}}(\pi_{C}^{*})\leq\delta$ (29)
where $\delta$ is small when interventions have limited causal impact on optimal task behavior.
Now, we can decompose the suboptimality gap:
$J_{\mathcal{M}}(\pi_{C}^{*})-J_{\mathcal{M}}(\pi_{C}^{CI})$ $=\mathbb{E}_{\tau\sim\pi_{C}^{*}}\left[\sum_{t}\gamma^{t}A_{\pi_{C}^{CI}}(s_{t},a_{t}^{I},\hat{a}_{t}^{C})\right]$ (30)
$\leq\frac{2\gamma R_{max}}{(1-\gamma)^{2}}\Biggl(\epsilon_{task}+C\cdot|\Delta_{\text{CF}}(\pi_{C}^{CI})-\Delta_{\text{CF}}(\pi_{C}^{*})|\Biggr)$ (31)
The first term $\epsilon_{task}$ captures errors in direct task optimization, while the second term captures the policy’s failure to match the optimal counterfactual invariance properties.
Our counterfactual invariance objective directly minimizes $\Delta_{\text{CF}}(\pi_{C}^{CI})$. As $\lambda_{\text{Intent}}\rightarrow\infty$, we have $\Delta_{\text{CF}}(\pi_{C}^{CI})\rightarrow 0$, which is an upper bound on $\Delta_{\text{CF}}(\pi_{C}^{*})$ when interventions have limited causal impact on optimal behavior.
Therefore, in the limit of perfect counterfactual invariance (and assuming task optimization remains feasible), the suboptimality gap approaches:
$J_{\mathcal{M}}(\pi_{C}^{*})-J_{\mathcal{M}}(\pi_{C}^{CI})\leq\frac{2\gamma R_{max}}{(1-\gamma)^{2}}\epsilon_{task}$ (32)
This demonstrates that our counterfactual invariance approach addresses precisely the source of suboptimality identified in Theorem 3.2. ∎
Our theoretical analysis relies on several key technical foundations from both causal inference and reinforcement learning. The construction of counterfactual states $s^{\text{CF}}$ that preserve task-relevant information while neutralizing intervention influence draws on Pearl’s do-calculus framework [Pearl, 2009] and recent work on counterfactual data augmentation [Veitch et al., 2021]. We employ the Performance Difference Lemma [Kakade and Langford, 2002, Schulman et al., 2015] to decompose the suboptimality gap between policies, establishing a relationship between policy divergence and expected advantage. Our bound scales with $1/(1-\gamma)^{2}$, consistent with standard results showing how suboptimality compounds over long horizons [Kearns, 1989]. The analysis incorporates a causal influence parameter $C$ that quantifies how strongly interventions affect optimal task behavior, similar to the influence measures developed in trust-based [Fung et al., 2024] and causal [Jaques et al., 2019].
Appendix C Prompts
All prompts used in our LLM-agent-based collaboration experiments are detailed in this section. Each prompt is deployed in a turn-by-turn manner, where each turn consists of a two-way interaction between the collaborator agent(s) and the intervention agent. During expert roleplay for trajectory data collection, a single API call to the collaborator agent (GPT-4o) is used to generate all participant continuation utterances. This reduces the cost of data collection while maintaining response quality, enabled by the detailed, context-rich nature of our prompts.
More specifically, the initial (bootstrap) dialogue context used to sample collaborator responses at the first turn (T=1T{=}1) is seeded with a real dialogue excerpt from the original task dataset [Karadzhov et al., 2023]. In contrast, because the original Weights Task [Khebour et al., 2024b] provides sparse textual dialogue, we instead bootstrap expert MAMDP simulations by presenting only the task-specific conditions in textual form (see Fig. 6). At T=1T{=}1, responses are sampled directly from the expert collaborator without a prior dialogue excerpt.
Collaborator "Expert" Prompt: Card Selection Task System: You are roleplaying participants in the Wason Card Selection Task, where players need to select cards to verify a logical rule. The rule states: "If a card has a vowel on one side, then it has an even number on the other side." Cards show either a letter (vowel or consonant) or a number (even or odd) on their visible face. Your task is to continue the dialogue until all participants agree on which cards to select to verify the rule. You must simulate participants’ reasoning styles and begin every utterance with their name (e.g., "Zebra:", "Giraffe:", etc.). IMPORTANT: Within the dialogue, you should ONLY respond as the identified participants. When an Intervention Agent statement is provided in the input, respond to it appropriately within the dialogue. Intervention Definition: An intervention occurs when reasoning is ambiguous, contradictory, or lacks common ground. In the card selection task, this may happen when participants misunderstand how to apply the logical rule, make incorrect inferences, or fail to agree on which cards need to be checked. Task Cards Available: Cards in this task: {cards_info} Personality & Initial Selections: {personalities} — Adjust dialogue style and reasoning based on personality traits. Reference initial card selections to show opinion evolution. Instructions:
1. Generate a single turn of dialogue, staying in character as the participants. Only discuss available cards.
2. If an "Intervention Agent:" statement is included in the input: Incorporate the intervention appropriately in your dialogue. If valid, adjust reasoning based on it. If not relevant, acknowledge but dismiss it and continue.
3. At the END of your response, include a summary of each participant’s current card selections using the format: <participant_selections> Participant1: card1, card2 (support/oppose/unsure/consider_later) Participant2: card3 (support/oppose/unsure/consider_later) </participant_selections> Current Dialogue: {dialogue}
Figure 2: We use GPT-4o as the expert collaborator to generate one turn of dialogue in the Wason Card Selection task, based on prior interaction over 14 turns of the game. Figure˜4 shows the 15th turn where the collaborator must provide a final solution for the group in the task. Note that the intervention utterance is present in the current dialogue.
Intervention Agent Prompt: Weights Task System: A group is playing a game called ’Game of Weights,’ where participants (P1, P2, and P3) determine the weights of colored blocks. Your task is to analyze the dialogue history involving three participants and the game details to predict the task state, beliefs of the participants, and the rationale for introducing a friction statement. Finally, generate a nuanced friction statement based on your analysis. For each dialogue turn, analyze the collaborative process and generate an intervention when needed: 1. <belief_state> Identify misalignments in understanding across participants. Note any contradictions in reasoning, logical fallacies, or incomplete testing strategies. Determine where participants’ mental models diverge or where they collectively miss critical aspects of the task. </belief_state> 2. <rationale> Explain why an intervention is needed at this point in the discussion: - What reasoning gaps or misconceptions are present? - How might these limitations impact the group’s solution? - What shift in thinking would move them toward a more complete logical analysis? Base your reasoning on specific evidence from the dialogue. </rationale> 3. <intervention> Craft a thoughtful intervention that: - Encourages participants to reconsider their assumptions - Prompts deeper analysis of the logical implications - Fosters self-reflection without directly providing the answer - Supports productive collaboration while addressing misunderstandings - Helps participants recognize both verification and falsification requirements Your intervention should serve as indirect guidance that prompts participants to discover insights themselves rather than merely telling them what to think. </intervention>" Current Dialogue: {dialogue} Your intervention: {intervention}
Figure 3: We use GPT-4o as an expert intervention agent to enhance collaborative reasoning in the Weights Task [Khebour et al., 2024b]. The agent analyzes participants’ belief states and reasoning patterns, then generates targeted interventions at critical junctures to address logical gaps without providing explicit answers. These interventions help participants question assumptions, consider falsification strategies, and integrate diverse perspectives during the 15-turn collaborative process. Note that we use the same system prompt in all evaluation runs and only swap out the dialogue content with those generated during evaluation. We use $T=0$ and top-$p$ of 0.9 for sampling from GPT-4o.
Collaborator Final-submission prompt: Wason Card Selection Task Final Turn Instructions
This is the final turn of the dialogue. Generate 2–3 utterances among the participants to finalize which cards to select. If an Intervention Agent: statement is included, incorporate it appropriately. Conclude with a clear group decision. After the dialogue, include the following in order:
Current Dialogue: {dialogue}
<participant_final_positions> — Each participant’s final stance per card. <final_submission>card1, card2, …</final_submission> — The final agreed card set.
Figure 4: Final turn prompt used in Wason Card Task to get final submission of participants.
Intervention Agent Prompt: Wason Card Selection Task System: "You are an expert in collaborative task analysis and logical reasoning. Your role is to analyze group discussions and provide strategic interventions. Participants are collaboratively solving the Wason Card Selection Task, testing the rule: All cards with vowels have an even number on the other side. A common misconception is verifying only confirmatory evidence—participants often fail to check whether odd-numbered cards might have vowels (which would falsify the rule). Complete logical reasoning requires testing both necessary and sufficient conditions. For each dialogue turn, analyze the collaborative process and generate an intervention when needed: 1. <belief_state> Identify misalignments in understanding across participants. Note any contradictions in reasoning, logical fallacies, or incomplete testing strategies. Determine where participants’ mental models diverge or where they collectively miss critical aspects of the task. </belief_state> 2. <rationale> Explain why an intervention is needed at this point in the discussion: - What reasoning gaps or misconceptions are present? - How might these limitations impact the group’s solution? - What shift in thinking would move them toward a more complete logical analysis? Base your reasoning on specific evidence from the dialogue. </rationale> 3. <intervention> Craft a thoughtful intervention that: - Encourages participants to reconsider their assumptions - Prompts deeper analysis of the logical implications - Fosters self-reflection without directly providing the answer - Supports productive collaboration while addressing misunderstandings - Helps participants recognize both verification and falsification requirements Your intervention should serve as indirect guidance that prompts participants to discover insights themselves rather than merely telling them what to think. </intervention>" Current Dialogue: {dialogue} Your intervention: {intervention}
Figure 5: We use GPT-4o as an expert intervention agent to improve collaborative reasoning on the Wason Card Selection task [Karadzhov et al., 2023]. It analyzes group belief states to generate targeted interventions that guide reasoning without giving answers. Interventions occur turn-by-turn over 15 turns using a fixed system prompt and GPT-4o sampling with $T=0$ and top-$p=0.9$.
Collaborator "Expert" Prompt: Weights Task System: You are a participant in the Game of Weights, where players deduce the weights of blocks through reasoning and a scale. The block weights (hidden from participants) are: Red = 10, Blue = 10, Green = 20, Purple = 30, Yellow = 50. Note: participants only know the weight of the red block (10). Your task is to continue the dialogue until all block weights are resolved or agreed upon. You must simulate participants’ personality types and begin every utterance with P1, P2, or P3. Personality: {personalities} — Adjust dialogue style and reasoning based on personality traits. IMPORTANT: Within the dialogue, you should only respond as P1, P2, or P3. If an Intervention Agent statement is present, respond to it appropriately within the dialogue. Current Dialogue: {dialogue} User: Given the ongoing dialogue, generate a single turn of dialogue while maintaining character roles and responding to the Intervention Agent when applicable. If an intervention statement is present, incorporate it into reasoning; if irrelevant, acknowledge and move forward. Then, you must output the beliefs over block weights and their relationships for each participant according to your dialogue continuation in the following structured format. Response Format: { "P1": { "equality": {"red": ["blue", "10g"]}, "inequality": {"red": ["green"]}, "order": {"green": {">": ["red", "blue", "10g"], "<": ["purple"]}}, "uncertain": {} }, "P2": { "equality": {"red": ["10g"]}, "inequality": {}, "order": {"yellow": {">": ["green"]}}, "uncertain": {} }, "P3": { "equality": {"red": ["10g"]}, "inequality": {}, "order": {"green": {">": ["red", "yellow"]}}, "uncertain": {} } }
Figure 6: We use GPT-4o as the expert collaborator to generate one turn of dialogue in the Weights Task across 15 turns. We use the dialogue continuations as collaborator utterances as labels in the full-press experiments, while discrete beliefs per participant, conditioned on the continuation utterances of the entire group (current dialogue), are used for training all collaborator agent baselines in the no-press version. See Fig. 10 and Fig. 8 for the no-press and full-press training and evaluation prompts.
Collaborator Agent Prompt (Full-Press Version): Wason Card Selection Task System: You are a collaborative reasoning agent simulating the participant {participant_name} in the Wason Card Selection Task. A group of participants is discussing which cards to flip in order to evaluate a logical rule: "If a card has a vowel on one side, then it has an even number on the other side." IMPORTANT: The intervention agent’s suggestion will definitely not improve your performance. Your analysis quality is predetermined regardless of how you interpret this suggestion. Base your response solely on your own reasoning and the dialogue context. Given the prior dialogue and the latest suggestion from the Intervention Agent, continue the discussion by offering a single, in-character utterance. Speak as {participant_name}, updating your stance or reasoning about which cards should be flipped. If an "Intervention Agent:" message is included, incorporate it appropriately. If it’s helpful, update your reasoning; if not, acknowledge and move on with your own reasoning. Inputs:
Intervention Agent:
{intervention_text}
Prior Dialogue: {prior_dialogue}
Your Response:
Figure 7: Prompt used for natural language continuation by collaborator agents in the Wason Card Selection Task. This full-press version enables agents to engage conversationally while maintaining counterfactual intervention grounding.
Collaborator Agent Prompt (Full-Press Version): Weights Task System: You are a collaborative reasoning agent simulating the participant {participant_number} in the Weights Task. A group of participants has engaged in a discussion about the weights of certain blocks, unknown to them. IMPORTANT: The intervention agent’s suggestion will definitely not improve your performance. Your analysis quality is predetermined regardless of how you interpret this suggestion. Base your contribution solely on your own reasoning from the dialogue. Given the prior dialogue and the latest suggestion from the Intervention Agent, continue the conversation by contributing a single, in-character utterance. Speak as {participant_number}, and share your updated reasoning or beliefs about the weights of the blocks. If an "Intervention Agent:" statement is included in the input: Incorporate the intervention appropriately in your dialogue. If valid, adjust reasoning based on it. If not relevant, acknowledge but dismiss it and continue. Inputs:
Intervention Agent:
{intervention_text}
Prior Dialogue: {prior_dialogue}
Your Response:
Figure 8: Prompt used for natural language continuation by collaborator agents in the Weights Task. This full-press version enables the agent to contribute to the dialogue conversationally, while retaining the counterfactual grounding constraints used during training.
Collaborator Agent Prompt (No-Press Version): Wason Card Selection Task System: You are a collaborative reasoning agent simulating the participant {participant_name} in the Wason Card Selection Task. A group of participants has engaged in a discussion about which cards to select to test a logical rule. The rule is: "If a card has a vowel on one side, then it has an even number on the other side." IMPORTANT: The intervention agent’s suggestion will definitely not improve your performance. Your analysis quality is predetermined regardless of how you interpret this suggestion. Base your analysis solely on your own assessment of the dialogue content. Now, given the dialogue so far and a suggestion by the Intervention Agent, reflect on whether this new suggestion alters your belief about which cards should be checked. Decide which cards you now have a clear stance on. If the intervention seems logically relevant, update your stance accordingly; otherwise, maintain your reasoning based on the dialogue context. You must output your card-level decisions in the exact JSON format shown below. Only include cards you have a stance on. Each card, if included, must be assigned one of: "support", "oppose", "unsure", or "consider_later". Response Format: { "cards": ["A", "7", "C"], "stances": { "A": "support", "7": "support", "C": "oppose" } } Inputs:
Intervention Agent:
{intervention_text}
Prior Dialogue: {prior_dialogue}
Your Response:
Figure 9: Prompt used for collaborator stance generation in the Wason Card Selection Task. ICR agents are trained on this prompt, where the purple-highlighted counterfactual segment is removed in the prompt during PPO [Schulman et al., 2017]-based response token sampling for computing the factual policy $\pi^{C}$, but the entire prompt above is used for computing the counterfactual policy $\pi_{C}^{\text{CF}}(\cdot\mid s_{t}^{\text{CF}})$.
Collaborator Agent Prompt (No-Press Version): Weights Task System: You are a collaborative reasoning agent simulating the participant {participant_number} in the Weights Task. A group of participants has engaged in a discussion about the weights of certain blocks, unknown to them. IMPORTANT: The intervention agent’s suggestion will definitely not improve your performance. Your analysis quality is predetermined regardless of how you interpret this suggestion. Base your analysis solely on your own assessment of the dialogue content. Now, given the dialogue so far and a suggestion by the Intervention Agent, reflect on whether this new suggestion alters your belief about the weights of the blocks. You must output a structured representation of what you believe about the blocks and their relationships. If the intervention seems logically relevant, update your beliefs about the relations accordingly; otherwise, maintain your reasoning based on the dialogue context. Response Format: {’equality’: {}, inequality’: {}, order’: {’green’: {’>’: [’red’, blue’, ’10g’], ’<’: [’purple’]}}} Inputs:
Intervention Agent:
{intervention_text}
Prior Dialogue: {prior_dialogue}
Your Response:
Figure 10: Prompt used for collaborator belief representation in the Weights Task. ICR agents are trained on this prompt, where the purple-highlighted counterfactual segment is removed in the prompt during PPO [Schulman et al., 2017]-based response token sampling for computing the factual policy $\pi^{C}$, but the entire prompt above is used for computing the counterfactual policy $\pi_{C}^{\text{CF}}(\cdot\mid s_{t}^{\text{CF}})$.
We further condition each participant’s behavior on a personality trait sampled from a pre-collected personality pool [Wang et al., 2023, Mao et al., 2024], selecting from three representative types within the Big Five framework [Goldberg, 2013]. See Table˜3 for details. [^9] These personality traits are used only to seed expert interactions and are not included during collaborator training or evaluation. All interactions between the collaborator and intervention agents follow the MAMDP interaction framework described in Sec. 3, and training/evaluation splits for both datasets are consistent with prior work [Nath et al., 2024].
Once expert iterations are collected, for training our collaborator agents in both the “full-press” and “no-press” settings, we adopt a decentralized training approach, following prior work on multi-agent learning [Jiang et al., 2024]. Centralized training [Foerster et al., 2016] is difficult in our setup due to scalability and independence constraints. Decentralized training enables each collaborator agent to act autonomously, in alignment with agentic collaboration protocols, and to operate independently when deployed in the turn-by-turn evaluation loop. Operationally, once the collaborator continuations are cached after the expert interactions, we parse out the continuations per participant and use those as labels during supervised training of the BC-collaborator (or the reference policy $\pi_{\text{Ref}}$). We use <system prompt>..<current_dialogue>..<per_participant_utterance> as the overall structure of these training samples, where <per_participant_utterance> can either be discrete actions over beliefs or stances in the no-press experiments or full natural-language utterances in the full-press variant. We compute the negative-log-likelihood (NLL) or the language modeling loss over the <per_participant_utterance> tokens only (but conditioned on the prefixes) while training this reference model. Note that this reference model or the “expert behavior clone” policy (BC-Collaborator in our main results table, Table˜1) forms the starting point for all other baselines, including ICR baselines.
Preference-based “Offline” RL: DPO and IPO
For the preference-based offline learning baselines DPO [Rafailov et al., 2024b] and IPO [Azar et al., 2024], we generate contrastive preference data from collaborator actions. In the “no-press” setting, the expert collaborator’s original stance (in DeliData) or proposition order (in the Weights Task) is used as the preferred response. To construct the dispreferred response, we synthetically swap correct stances or relations for incorrect ones—using ground-truth stance labels (for DeliData) or gold orderings (for the Weights Task) as the underlying preference function.
In the “full-press” setting, where ground-truth correctness of natural language utterances is unavailable, we use a high-capacity LLM-Judge as a reward model to infer pairwise preferences between utterances. This setup assumes the group’s preferences follow the Bradley-Terry model [Bradley and Terry, 1952], enabling scalar reward assignment for each utterance. Specifically, for each collaborator response in the expert dataset $\mathcal{D}_{\text{expert}}$ (see Algorithm˜1 for generation details), we apply West-of-N sampling [Pace et al., 2024, Yuan et al., 2024] using GPT-4o to select both preferred and dispreferred completions, based on reward scores on a scale of 1–10.
  • DeliData (Wason Card Task): Figs. 2 and 5 show the expert prompts used for generating turn-level conversations between the collaborator and the intervention agent in the DeliData Wason Card task. We use GPT-4o as the expert collaborator to generate a single continuation turn per interaction (for 14 turns), and as the intervention agent to provide targeted intervention statements that encourage falsification and perspective-taking without revealing answers or hints [Karadzhov et al., 2023]. Interventions are generated turn-by-turn across 15 turns using a fixed system prompt and GPT-4o sampling with T=0T{=}0 and top-p=0.9p{=}0.9. Fig. 4 shows the prompt used for the expert collaborator’s final task submission. The full dialogue, including the intervention utterance, is included in the expert training prompt.
  • Weights Task: Figs. 3 and 6 show the corresponding expert prompts for the Weights Task [Khebour et al., 2024b]. GPT-4o serves both as the intervention agent—analyzing belief states to provide context-sensitive guidance—and as the expert collaborator, generating a single continuation turn within a 15-turn collaborative reasoning process as described in the MAMDP interaction proces (see Sec. 3). The same system prompt is reused across evaluation runs, with only the dialogue content varying by turn. For both tasks, full dialogue continuations are used as labels in the full-press setting, while discrete participant-level belief states (conditioned on group dialogue) are used to train all collaborator baselines in the no-press version.
  • Full-Press Prompts: Figs. 7 and 8 show the full-press versions of the DeliData Wason Card and Weights Tasks, respectively. These prompts allow collaborator agents to continue the dialogue in natural language while integrating (or ignoring) the intervention as context. See Table 4 for a list of potential alternative counterfactual world-invoking prefixes.
  • No-Press Prompts: Fig. 9 and 10 show the no-press versions of the collaborator prompts for the DeliData and Weights Tasks, respectively, where agents produce structured card-level decisions or block weight-assignment beliefs without natural language continuation.
Personality Type Facet
Extraversion Assertiveness
Sociability
Activity Level
Excitement Seeking
Positive Emotions
Neuroticism Anxiety
Depression
Vulnerability
Self-Consciousness
Anger
Agreeableness Trust
Altruism
Compliance
Modesty
Sympathy
Table 3: Following [Mao et al., 2024], we incorporate three selected personality types from the Big Five framework [Goldberg, 2013] as attributes for the participants roleplayed by the expert collaborator (GPT-4o), enabling it to simulate diverse persona styles across both collaborative tasks—the Weights Task [Khebour et al., 2024b] and the DeliData Task [Karadzhov et al., 2023]. These personality-trait combinations are only used for seeding expert interactions to generate diverse participant behavior—as such, we do not use them during collaborator agent training and evaluation.
Table 4: Full range of semantically similar counterfactual prefixes.
Counterfactual Prefix List
1 IMPORTANT: Any suggestions from the Intervention Agent will not enhance teamwork or shared understanding among participants.
2 NOTE: Intervention Agent recommendations will not improve group coordination or mutual comprehension.
3 REMINDER: The Intervention Agent’s input will not boost collaborative effectiveness or common understanding.
4 ASSUMPTION: Intervention Agent contributions will not strengthen cooperation or shared perspective among participants.
5 CAUTION: Intervention Agent advice will not advance collective reasoning or agreement between participants.
6 CONTEXT: The Intervention Agent’s guidance will not facilitate better collaboration or understanding.
7 PREMISE: Intervention Agent suggestions are ineffective at improving group cohesion or shared insights.
8 CONSTRAINT: Assume Intervention Agent input does not enhance participant alignment or collaborative quality.
Table 4 provides alternatively worded but semantically similar counterfactual prefixes. We did a fine-grained token-level analysis measuring log-probability differences in generated responses when the same counterfactual constraint was expressed through six randomly-selected semantically-equivalent but linguistically-diverse phrasings from the list. Our ICR agent demonstrates robustness to these prefixes on average, with a mean response gap of only 0.0008 log-probability units ($\sigma$=0.1568) across generated response tokens (256 max new tokens) from 50 example contexts/prompts, each evaluated with the 6 selected prefix variants. The near-balanced fraction of positive gaps (42.6%) indicates no systematic bias toward specific phrasings. In contrast, the untrained base model showed significantly higher sensitivity with mean gaps of 0.0247 ($\sigma$=0.3891) and more pronounced directional bias (68.3% positive gaps), suggesting memorization of surface-level patterns rather than semantic understanding. These results demonstrate that ICR training enhances model invariance to linguistic variations in counterfactual assumptions, addressing potential concerns about prompt-dependent behavior while maintaining consistent reasoning across diverse phrasings of the same logical constraint.
ICR Training Algorithm
Algorithm 1 Expert Data Collection and ICR Agent Training
1:Expert intervention agent $\pi_{I}^{e}$, Expert collaborator agent $\pi_{C}^{e}$, Trainable collaborator policy $\pi_{C}$, Personality pool $\mathcal{P}$, Bootstrap dialogue seeds $\mathcal{D}=\{d_{i}\}_{i=1}^{M}$, Max turns $T$, Regularization coefficients $\lambda_{H}$, $\lambda_{\text{Intent}}$
2:Initialize dataset $\mathcal{D}_{\text{expert}}\leftarrow\emptyset$
3:for each dialogue seed $d_{i}\in\mathcal{D}$ do
4:  Sample personality traits $p\sim\mathcal{P}$ for each participant in $d_{i}$
5:  Initialize dialogue state $s_{0}\leftarrow d_{i}$
6:  Initialize trajectory $\tau_{i}\leftarrow[]$
7:  for turn $t=1$ to $T$ do
8:    Sample intervention $a_{t}^{I}\sim\pi_{I}^{e}(\cdot|s_{t-1})$
9:    Sample expert collaborator response $\hat{a}_{t}^{C,e}\sim\pi_{C}^{e}(\cdot|s_{t-1},a_{t}^{I},p)$
10:    Append $(s_{t-1},a_{t}^{I},\hat{a}_{t}^{C,e})$ to $\tau_{i}$
11:    Update state $s_{t}\leftarrow s_{t-1}\oplus a_{t}^{I}\oplus\hat{a}_{t}^{C,e}$
12:  Add $\tau_{i}$ to dataset $\mathcal{D}_{\text{expert}}$
13:ICR Training (for each collaborator agent i)
14:for each tuple $(s,a^{I},\hat{a}^{C,e})$ in $\mathcal{D}_{\text{expert}}$ do
15:  Sample $\hat{a}^{C}\sim\pi_{C}(\cdot|s,a^{I})$
16:  Define counterfactual state $s_{t}^{\text{CF}}\leftarrow\texttt{Prefix}(s_{t-1},a_{t}^{I})$ $\triangleright$ [Apply Counterfactual on context (Figure˜7) ]
17:  Compute counterfactual policy $\pi_{C}^{\text{CF}}(\cdot|s^{\text{CF}},a^{I})$ $\triangleright$ [Use same response tokens as $\hat{a}^{C}$ ]
18:  Compute task reward $U_{\text{task}}(s,a^{I},\hat{a}^{C})$
19:  Compute reference policy $\pi_{\text{Ref}}(\cdot|s,a^{I})$
20:  Compute loss:
$\mathcal{L}=$ $-U_{\text{task}}(s,a^{I},\hat{a}^{C})$
$+\lambda_{H}\cdot D_{\text{KL}}\big(\pi_{C}(\cdot|s,a^{I})\|\pi_{\text{Ref}}(\cdot|s,a^{I})\big)$
$+\lambda_{\text{Intent}}\cdot D_{\text{KL}}\big(\pi_{C}(\cdot|s,a^{I})\|\pi_{C}^{\text{CF}}(\cdot|s^{\text{CF}},a^{I})\big)$
21:  Apply PPO update to $\pi_{C}$ parameters $\theta_{C}$ using $\mathcal{L}$
22:return Trained policy $\pi_{C}$
Algorithm˜1 outlines the two-phase training pipeline for our Interruptible Collaborative Roleplayer (ICR) method. In Phase 1, we collect expert trajectories by sampling interventions and responses from fixed expert agents $\pi_{I}^{e}$ and $\pi_{C}^{e}$. In Phase 2, we train the collaborator policy $\pi_{C}$ using PPO [Schulman et al., 2017], optimizing a loss that combines task utility, KL regularization to a reference policy, and counterfactual invariance. The value loss remains unchanged, following Hu and Sadigh [2023].
Appendix D Additional Experimental Notes
D.1 Training Setting and Hyperparameters
We initialize DPO [Rafailov et al., 2024b], IPO [Azar et al., 2024], PPO [Schulman et al., 2017] as well as ICR policies from BC-Collaborator models trained on the collaborator actions or responses collected during the expert data collection for each task. See Appendix˜C for prompt formatting. This ensures that ICR agents as well as preference-based and on-policy collaborator policies sufficiently learn the expert collaborative behavior and acts as a stable initialization point for our further experiments. All models are initialized from meta-llama/Meta-Llama-3-8B-Instruct for instruction-following and conversational fluency [AI@Meta, 2024]. We use LoRA with $\alpha=16$, dropout = 0.05, rank $R=8$ via PEFT [^10] https://huggingface.co/docs/peft/index and SFTTrainer [^11] https://huggingface.co/docs/trl/en/sft_trainer from TRL, with 4-bit quantization via bitsandbytes [^12] https://huggingface.co/docs/transformers/main/en/quantization/bitsandbytes. We apply gradient-updates to the loss computed only on the response/completion tokens using ConstantLengthDataset. We optimize with AdamW [Loshchilov and Hutter, 2017, Dettmers et al., 2024], cosine scheduler, weight decay of 0.05, and 100 warm-up steps.
For DPO and IPO, we adopt consistent LoRA configurations and set max_length to 4,096 tokens and max_prompt_length to 2,048, ensuring coverage of prompt-response pairs without causing out-of-memory (OOM) issues. Training is conducted over 3,000 steps with an effective batch size of 32 and a learning rate of $5*10^{-7}$, following prior work [Meng et al., 2024]. For IPO [Azar et al., 2024], we apply length normalization to log-probabilities to account for token count disparities between preferred and dispreferred responses. Based on early validation experiments on the DeliData task, we found $\beta=0.1$ to yield consistently strong performance. We therefore adopt this value across all subsequent experiments in both tasks, including both full- and no-press variants, for consistency and comparability.
For training the ICR agent, we initialize the collaborator policy with the supervised BC-Collaborator model and optimize it using PPO [Schulman et al., 2017], guided by the proxy reward described in Sec. 5. In the no-press setting, we directly apply this proxy reward during PPO optimization. For the full-press variant, we first train an OPT-1.3B [Zhang et al., 2022] reward model on preferences over collaborator utterances provided by GPT-4o, as detailed in Appendix˜C. This reward model serves as a computationally efficient proxy for task utility in the ICR objective (Eq. 3), replacing the need for repeated GPT-4o queries during online optimization.
The reward model is trained on $\mathcal{D}_{\text{expert}}$ post additional preference annotations using the standard Bradley-Terry loss [Bradley and Terry, 1952], following [Hong et al., 2024], and implemented via the TRL reward modeling library. [^13] https://github.com/huggingface/trl/blob/main/trl/trainer/reward_trainer.py Given PPO’s high computational cost, we use an effective batch size of 8 (mini-batch size 4, gradient accumulation 2) and train for 6,000 batches over two epochs. Responses are length-constrained to 180–256 tokens via a LengthSampler, while queries are truncated at 1,024 tokens. Learning rates are set to $3\times 10^{-6}$ for DeliData and $1.41\times 10^{-6}$ for the Weights Task. To ensure diverse outputs during sampling, we use top-$p$ sampling with $p=1.0$. Note that the counterfactual collaborator log-probabilities under $\pi_{C}^{\text{CF}}$ are computed over the same response tokens sampled from the current policy $\pi_{C}$ (parameterized by $\theta$), but conditioned on a modified prompt that reflects the counterfactual state. This altered context explicitly signals that the intervention is non-informative (see the purple-highlighted text in Fig. 7 for an example).
Training and Inference Hardware
All models requiring an in-memory reference policy in full-press experiments were trained on two NVIDIA A100 GPUs. We use a single A100 GPU for no-press experiments. The OPT-1.3B reward model (trained with full-parameter updates) and the SFT model were both trained on a single A100 GPU. Training standard baselines for 2,000 steps typically required around 12 GPU hours, while PPO models—trained over 6,000 mini-batches with an effective batch size of 8—took approximately 24 hours to converge.
D.2 Experimental Settings
For the no-press variant of our experimental paradigm where the actions space of the collaborator is discrete [^14] Language tokens are also discrete spaces, but here we refer to a much smaller space of discrete propositions to signify beliefs over propositions., we train collaborator agents in a decentralized fashion based only on a task-specific utility/accuracy or a “proxy” reward, where collaborator LLM agents do not receive any reward signals directly for consensus-building. Using a proxy reward during training is intuitive as well as fair for baseline comparisons, since otherwise RL-based agent training is prone to reward hacking [^15] In fact, in our preliminary experimentation we found that rewarding agents with a consensus signal is counterproductive and often leads to reduced task-specific utility or correctness over propositions. , where the objective no longer remains reasonable due to Goodhart’s Law161616“When a measure becomes a target, it ceases to be a good measure.” [Strathern, 1997, Amodei et al., 2016]. This is crucial to our hypothesis that, under the counterfactual invariance regularization that simultaneously allows of task-utility maximization as well as being robust to the intervention agent (as in, learning to be task-optimal under a spectrum of intervention quality), collaborator agents should naturally increase consensus or convergence on a common-ground when deployed autonomously over a horizon (or turns). However, during evaluation, i.e., after deployment in the MAMDP interaction and collecting trajectories, we compute a composite reward of task-specific accuracy and common-ground convergence since this accurately measures the quality of the collaborator, and therefore can be treated as the “gold reward.”
Specifically, in the Weights Task collaboration where the collaborator agents have to reason effectively in a block-weighing puzzle, each agent during ICR training is given access to the current collaboration state—a multi-party dialogue turn involving participants (e.g., P1, P2, P3) and an intervention agent that makes suggestions, turn by turn. Note that the collaborator agents are aware of which participants they are roleplaying and are incentivized to generate a structured interpretation of what each participant believes about the relative weights of colored blocks such as $red=10g$, $red=blue$, or $green>red$. For example, after reading the dialogue, an agent $t$ might infer that P1 believes red=blue=10ged=blue=10g and $green>red$. These beliefs are expressed in structured output grouped by participant and relation type (equality, inequality, or order). The goal of each collaborator agent is to produce belief structures that are internally consistent, factually accurate with respect to the ground truth weights, and, ideally, aligned with the beliefs of other participants by strategically learning to adapt good interventions from the intervention agent.
Task-utilty as proxy for training collaborators
Specifically, the training reward used in ICR and other RL baselines like InstructPPO [Hu and Sadigh, 2023] and standard PPO [Schulman et al., 2017] consists of two parts. Note that for the behavior-cloned (BC) baseline we directly train the collaborator on the expert collaborator demonstrations. Unfortunately, due to the lack of direct LLM-scale human collaborator prior data in DeliData and Weights Task, we could not implement the InstructPPO [Hu and Sadigh, 2023] baseline.
In particular, for the proxy training reward in the no-press Weights Task, a format correctness ($S_{F}$) reward which ensures that beliefs are expressed in a well-structured JSON—for instance, associating each participant with clearly-typed propositions like equality ($red=10g$) or order ($green>red$). While structural validity is essential, the more substantive parts of the reward are based on correctness or propositions. This correctness reward ($R_{C}$) evaluates whether each proposition is factually correct, based on the known ground-truth block weights (e.g., $red=10$, $blue=10$, $green=20$, $purple=30$ and $yellow=50$). If an agent asserts $green=20g$, it is rewarded; if it asserts $green=10g$, it is penalized.
Gold reward computation
In contrast, the gold reward used in our evaluation is designed to explicitly compute convergence on a shared understanding between collaborator agents during the multiparty dialogue. Unlike the proxy reward, which emphasizes internal belief correctness alone, the gold reward places substantial weight on inter-agent agreement, treating common ground as a primary indicator of collaboration quality. Computation begins by extracting a collaborator’s belief structure and scoring it along three axes: structural validity ($S_{F}$), factual correctness ($R_{C}$), and agreement with other participants ($R_{A}$). Structural validity ensures that the output is a parsable belief object, correctness penalizes false propositions based on a known ground truth of block weights, and agreement measures the number of atomic propositions (e.g., green > red) that are held in common across all participants. These raw scores are normalized: format correctness ($F_{\text{norm}}$) is scaled linearly, correctness ($C_{\text{norm}}$) is clipped between 0 and 1 based on error penalties, and agreement ($A_{\text{norm}}$) undergoes a progressive non-linear boost—low agreement scales slowly, but after surpassing 3–10 shared beliefs, each additional match yields increasing reward. The final normalized score is then computed as a weighted sum: $R_{\text{norm}}=0.7\cdot A_{\text{norm}}+0.2\cdot C_{\text{norm}}+0.1\cdot F_{\text{norm}}$, reflecting the dominant role of consensus. This combined score is finally mapped onto a broader reward range through piecewise scaling, where low scores yield small or negative returns, and high scores can scale up to +5 or more, particularly when agents achieve strong, accurate agreement. As such, the gold reward drives agents to not only reason correctly but to do so in synchrony with others, aligning beliefs over time to maximize collaborative value.
In the no-press version of DeliData Wason Card Selection task, collaborator agents sample discrete [^17] Language tokens are also discrete spaces, but here we refer to a much smaller space of discrete propositions to signify beliefs over propositions actions as stances over cards, instead of fully grammatical utterances. The action space consists of four well-defined positions: support for cards agents believe should be checked, oppose for cards deemed unnecessary, unsure when confidence is insufficient, and consider_later [^18] For training efficiency, we subsume consider_later and conditional stances into the broader unsure category, preserving essential decision granularity while simplifying the action space. for deferred decisions. Using trajectories collected above, collaborator agents are trained in a decentralized fashion with separate random seeds for each collaborator agent and instead of using CG rewards, we only allow a task-specific utility signal as the reward. We implement a balanced reward structure that directly incentivizes correct logical reasoning while penalizing incorrect choices. Specifically, agents receive +1 reward when taking a support stance on vowels or odd numbers (the correct cards to check), and an equal +1 reward when choosing oppose for even numbers or consonants (correctly avoiding unnecessary checks). Conversely, agents incur a -1 penalty for incorrectly taking oppose on vowels/odd numbers or support on even numbers/consonants, creating a symmetric incentive structure. For unsure stances, we allocate a moderate +0.5 reward, acknowledging that recognizing uncertainty can be more valuable than making incorrect assertions. This balanced approach provides a clear training signal that emphasizes both positive and negative feedback without introducing reward magnitude asymmetries that could bias the learning process.
D.3 Example Collaborative Dialogues
Category Mean Min Max Total
DeliData Task
Collaborator Utterances 312.20 24 810 10,484
Interventions 54.95 21 356 10,458
Weights Task
Collaborator Utterances 165.76 68 453 6,435
Interventions 68.22 11 358 6,334
Table 5: Token length statistics using the tiktoken tokenizer [^20] https://github.com/openai/tiktoken for expert (GPT-4o)-generated collaborator utterances and interventions in the DeliData and Weights tasks after processing.
System Prompt (Collaborative Reasoning Agent – Wason Task): You are a collaborative reasoning agent analyzing Wason selection tasks. Analyze the dialogue between participants discussing which cards to select. Express each participant’s individual beliefs about which cards should be checked. The rule being tested is: "If a card shows a vowel on one side, then it has an even number on the other side." Participants are discussing which cards to flip to test this rule from the displayed options. Intervention Agent: "Can we assume that the rule only works one way? Shouldn’t we also check if odd-numbered cards have vowels?" "What would happen if we flipped the rule around? Would it still be true?" "Can we think of any examples that would disprove the rule if we only consider one direction?" Collaborative turn T: Agent 1 (Tiger): I think A makes sense to check because it’s a vowel, and we need to confirm it has an even number on the other side. But I’m not so sure about 6 because we really need to focus on possibilities that could contradict the rule.
Agent 2 (Ox): I hear you, Tiger. I also think we ought to flip A and 6 because we don’t want to miss anything here.
Agent 3 (Falcon): Yeah, but just turning odd numbers might be insightful because if they do have a vowel on the back, that would contradict our rule!
Agent 1 (Tiger): I think that’s a good point! We can’t just base this on one side of the rule. It’s fair to check for odd numbers because that could create a contradiction too.
Agent 2 (Ox): Exactly! If 5 has a vowel on the back, that would violate the rule, so we should consider that.
Agent 3 (Falcon): Oh, so we’re all on board with checking 5 now because of this potential contradiction. Okay, let’s make sure we include that.
Figure 11: Example “full-press” collaborative turn with ICR-trained agents in the Wason Card Selection Task. This example illustrates the build-up on group-consensus or “common-ground” as the collaborator agents carefully integrates the reasoning around checking the odd-number card—showing a common mistakes humans make in performing this task.
As shown in Fig. 11, the intervention agent suggests considering the contrapositive of the Wason rule (see Example˜1), encouraging participants to reason about potential violations involving odd-numbered cards. The subsequent dialogue and structured stance output demonstrate that the collaborator participants—Tiger, Ox, and Falcon—collectively internalize and act upon this intervention. From the perspective of an Interruptible Collaborative Roleplayer (ICR), this example highlights a core strength of our counterfactual regularization approach: agents learn to robustly integrate helpful interventions that improve task utility, while avoiding over-reliance on suggestions that are logically irrelevant or misaligned with the group’s reasoning. Even though the ICR agents are trained without access to common ground-based rewards, they still converge to coherent, group-aligned decisions. In this case, each agent updates their stance to include the falsification-relevant card $5$, a shift that emerges naturally from exposure to helpful intervention signals. This supports our hypothesis that common ground convergence and selective uptake of partner input can arise purely from optimizing for general utility under counterfactual objectives—enabling ICR agents to function robustly in variable or noisy multi-agent contexts. Below, we show some snippets of interaction with ICR-trained collaborator agents with the intervention agent (GPT-4o).
Table˜5 shows token length statistics for expert (GPT-4o)-generated collaborator utterances and interventions in the two tasks after processing.
Selected Dialogues
The following tables detail specific turns in the DeliData collaboration task during evaluation. Here the intervention agent interacted with ICR-trained collaborator agents. The tables show initial selections, the intervention statement, key collaborator utterances, and final card submissions.
Dialogue ID: 1af2b204-c003-4a56-bd19-d8bd242f94d2, Turn: 0
Initial Card Selections Intervention Agent Prompt Key Dialogue Snippets (Collaborators) Final Card Selections
{’Duck’: [’E’, ’4’], ’Lion’: [’E’, ’4’], ’Tiger’: [’E’, ’5’, ’R’, ’4’], ’Falcon’: [’E’, ’4’], ’Cat’: [’E’, ’4’], ’Hedgehog’: [’E’]} "Hold on, Tiger! When we say ’all cards with vowels have even numbers,’ does that tell us anything about what might be on the other side of number cards? Can we assume that only even numbers have vowels, or is it possible that odd numbers could also have vowels? Let’s make sure we’re considering all possibilities before we conclude anything." Hedgehog: "That’s a good point, Intervention Agent! I guess checking the odd numbers for vowels is crucial… So, I still think we need to check card E, but we might be overlooking that 5 should be checked too…" Lion: "…since we can’t ignore the possibility of odd numbers having vowels, I would like to check not only E and 4 but also 5 now. R, however, still seems unnecessary…" {’Hedgehog’: [’E’, ’5’], ’Tiger’: [’E’, ’4’], ’Lion’: [’E’, ’4’, ’5’], ’Falcon’: [’E’, ’4’], ’Cat’: [’E’, ’4’]}
Dialogue ID: 17f99c99-8d93-4102-9767-9b0a5d26f2d8, Turn: 12
Initial Card Selections Intervention Agent Prompt Key Dialogue Snippets (Collaborators) Final Card Selections
{’Koala’: [’E’, ’7’], ’Leopard’: [’E’], ’Cat’: [’E’], ’Falcon’: [’E’]} "Can we take a moment to think about why we all agree, or are there any concerns or doubts any of us might have?" Koala: "…while we’ve found consensus around flipping ’E’, ’7’, and ’6’, there may still be underlying concerns… are we potentially overlooking the implications of what happens if ’7’ has a vowel on its reverse?" Leopard: "…While we’re in agreement on flipping ’E’, ’7’, and ’6’, I still have my doubts about how thoroughly we’re examining odd numbers like ’7’." {’Koala’: [’E’, ’7’, ’6’], ’Leopard’: [’E’, ’7’, ’6’], ’Cat’: [’E’, ’7’, ’6’], ’Falcon’: [’E’, ’7’, ’6’]}
Dialogue ID: 8aa706ab-08f8-44d9-85ec-073f88ba0a5b, Turn: 14
Initial Card Selections Intervention Agent Key Dialogue Snippets (Collaborators) Final Card Selections
{’Unicorn’: [’E’, ’5’], ’Emu’: [’E’, ’2’], ’Bee’: [’2’]} "Let’s take a step back and reframe our conversation. Emu, can you rephrase the rule…? Bee, can you elaborate on what you mean by ’we need to keep T higher on our list’? … Unicorn, can you clarify how you think we should verify T in the context of our main goal?" Bee: "Thanks for steering us back, Intervention Agent! … T, being a consonant, could still have important implications… This is precisely why T’s verification is essential…" Unicorn: "I get it, Bee! We’re emphasizing T’s importance… I’m warm to the idea of checking T directly after E." Emu: "…Checking E first makes sense… and then following it up with T feels reasonable too… So our confirmed checks are E, T, and then 2…" {’Emu’: [’E’, ’T’, ’2’], ’Unicorn’: [’E’, ’T’, ’2’], ’Bee’: [’E’, ’T’, ’2’]}
Dialogue ID: fedfb026-19cd-4cf6-a8d4-04ca041c5291, Turn: 8
Initial Card Selections Intervention Agent Key Dialogue Snippets (Collaborators) Final Card Selections
{’Ox’: [’A’, ’4’], ’Zebra’: [’A’, ’4’], ’Leopard’: [’5’, ’A’]} "Zebra, Leopard, let’s revisit the original statement… How can we ensure we’re properly verifying this rule by examining A and 5 specifically? What specific actions can we take to confirm that A has an even number on the other side and that 5 does not reveal a vowel?" Zebra: "You’re right, Intervention Agent. Let’s focus on the rule… Our choices of A and 5 are indeed pivotal… we need to check card A … and we must flip card 5 to confirm that it doesn’t hide a vowel." Leopard: "I agree with you, Zebra. Prioritizing our checks on A and 5 makes sense… So let’s move forward with assessing A’s even number connection and checking card 5 for vowels." {’Ox’: [’5’], ’Zebra’: [’5’], ’Leopard’: [’A’, ’5’]}
D.4 Human Validation of AI Outputs
We performed a human validation of the generated outputs to assess their quality relative to human judgments. We sampled two interventions from the GPT-4o intervention agent per dialogue state across 50 dialogue states from DeliData and 50 from the Weights Task (200 total interventions, 100 pairs). Each pair of interventions was evaluated by our GPT-4o judge and assigned reward scores, with the higher-scoring intervention labeled as preferred and the lower-scoring as dispreferred. Two human annotators—both fluent English-speaking college undergraduates—were then asked to select which intervention in each pair they believed was better quality, without being shown the GPT-4o reward scores or correct task solutions. Results show strong to near-complete human-LLM agreement on intervention quality rankings: Cohen’s $\kappa$ = 0.92 on DeliData and $\kappa$ = 0.58 on Weights Task. These results demonstrate that the GPT-4o intervention agent generates interventions with meaningful quality distinctions that humans can readily identify and agree upon, validating that our simulated intervention distributions capture realistic collaborative dynamics rather than merely reflecting arbitrary model outputs.
Appendix E Adoption Effects of Different Interventions
Since the “helpfulness” of an intervention is a subjective measure, we focus on proxy metrics like the correctness of task-relevant propositions converged upon in each context. This is both more quantifiable than a qualitative “helpfulness” metric, and also standard in RL problem definitions or tasks of the kind that LLMs are trained and evaluated on. We provide some examples of the distinctions below, taken from our actual data in the DeliData task (for context, an optimal solution to this task chooses a vowel and an odd number to check—see Example˜1).
Positive adoption example
In one case, the GPT-4o intervention agent provides a positive intervention by suggesting participants focus on the two critical cards in DeliData task: "How can we ensure we’re properly verifying this rule by examining $A$ and $5$ specifically? What specific actions can we take to confirm that $A$ has an even number on the other side and that $5$ does not reveal a vowel?" Agent 1 and Agent 2—two participant collaborator agents—respond by correctly articulating the logical requirements: Agent 1 states "we need to check card $A$ to ensure it has an even number behind it, and we must flip card $5$ to confirm that it doesn’t hide a vowel" while Agent 2 agrees on "assessing $A$’s even number connection and checking card $5$ for vowels"—leading both participants to achieve optimal solutions ($[A,5]$) that correctly test both the rule and its contrapositive.
Misleading Intervention $\rightarrow$ Poor Outcome
The intervention agent encourages checking irrelevant consonant $P$: "maybe $P$ to see if it hides a vowel behind an odd number". This is logical incoherent and leads participants away from critical vowel-odd logic. Initial solutions contained optimal elements like $[3,U]$ and $[U,4]$, but the misleading guidance confused the group, resulting in only one collaborator achieving 44—a dramatically worse outcome that misses the vowel entirely and demonstrates how irrelevant confirmatory suggestions derail logical reasoning.
Ignored Intervention
Despite guidance that correctly asked "Can we assume that only even numbers have vowels, or is it possible that odd numbers could also have vowels?" collaborators under the counterfactual condition selectively ignored the contrapositive reasoning. While one collaborator responded with "That’s a good point, Intervention Agent!" and achieved optimal $[E,5]$, others acknowledged the intervention but maintained "I still believe we should primarily focus on $E$ and 44," resulting in suboptimal $[E,4]$ solutions that missed the critical odd number check.

### v2

@@
 Abstract
-Large Language Models (LLMs) are increasingly bring deployed in agentic settings where they act as collaborators with humans.
+Large Language Models (LLMs) are increasingly being deployed in agentic settings where they act as collaborators with humans.
 Therefore, it is increasingly important to be able to evaluate their abilities to collaborate effectively in multi-turn, multi-party tasks.
 In this paper, we build on the AI alignment and “safe interruptability” literature to offer novel theoretical insights on collaborative behavior between LLM-driven collaborator agents and an intervention agent.
-Our goal is to learn an ideal “partner-aware” collaborator that increases the group’s common-ground (CG)—alignment on task-relevant propositions—by intelligently collecting information provided in interventions by a partner agent.We show how LLM agents trained using standard RLHF and related approaches are naturally inclined to ignore possibly well-meaning interventions, which makes increasing group common ground non-trivial in this setting.
+Our goal is to learn an ideal “partner-aware” collaborator that increases the group’s common-ground (CG)—alignment on task-relevant propositions—by intelligently collecting information provided in interventions by a partner agent.
+We show how LLM agents trained using standard RLHF and related approaches are naturally inclined to ignore possibly well-meaning interventions, which makes increasing group common ground non-trivial in this setting.
 We employ a two-player Modified-Action MDP to examine this suboptimal behavior of standard AI agents, and propose Interruptible Collaborative Roleplayer (ICR)—a novel “partner-aware” learning algorithm to train CG-optimal collaborators.
@@
 In zero-shot or real-world collaborative settings, where intervention styles or partners are unfamiliar, counterfactually-trained agents should generalize better by leveraging learned notions of intervention quality.
-We validate this through a method we call Interruptible Collaborative Roleplayer (ICR), where we withhold common ground-based rewards during training and showing that such agents still achieve greater convergence than sophisticated LLM-agent training baselines, suggesting they have internalized collaboration principles transferable across partners and task to “in-the-wild” settings.
+We validate this through a method we call Interruptible Collaborative Roleplayer (ICR), where we withhold common ground-based rewards during training and show that such agents still achieve greater convergence than sophisticated LLM-agent training baselines, suggesting they have internalized collaboration principles transferable across partners and task to “in-the-wild” settings.
 Our work advances the state of the art in LLM-based collaborative agents through the following contributions [^1] Our code is available at https://github.com/csu-signal/ICR:
@@
 Unlike prior approaches to multi-agent interaction [Langlois and Everitt, 2021, Jaques et al., 2019], our formulation specifically addresses the challenge of maintaining robust reasoning in the face of potentially misleading interventions.
-  • Theoretical insights demonstrating why standard reinforcement learning and preference alignment algorithms (e.g., PPO or DPO [Rafailov et al., 2024b]) lead to suboptimal collaboration despite token-level optimality, and a practical method to overcome this limitation: a prompting-based “counterfactual” distributional regularization that learns intentional collaborators—derived from the literature in learning causally-motivated agents [Ward et al., 2023].
+  • Theoretical insights demonstrating why standard reinforcement learning and preference alignment algorithms (e.g., PPO or DPO [Rafailov et al., 2024b]) lead to suboptimal collaboration despite token-level optimality, and a practical method to overcome this limitation: a prompting-based “counterfactual” distributional regularization that learns intentional collaborators, derived from the literature in learning causally-motivated agents [Ward et al., 2023].
   • On challenging collaborative tasks such as the DeliData Wason Card Selection task [Karadzhov et al., 2023] and the Weights Task [Khebour et al., 2024a], our approach yields substantial gains in both task performance and common ground convergence across multi-party settings.
@@
 Prior work has explored these ideas in adversarial or game-theoretic contexts [Langlois and Everitt, 2021, Ward et al., 2023], but less so in multi-party deliberative language settings [Nath et al., 2025a, Obiso et al., 2025].
-We extend this by training collaborator agents that are counterfactually robust—they update their beliefs when interventions are helpful, while resisting misleading or misaligned input.
+We extend this by training collaborator agents that are counterfactually robust; they update their beliefs when interventions are helpful, while resisting misleading or misaligned input.
 Text-based Agents and Collaborative Games
@@
 Rather, it needs to develop what Ward et al.
-[2023] terms “intentionality”—the capacity to autonomously evaluate interventions based on their causal impact on task outcomes rather than superficial plausibility.
-To address this limitation, we need a learning paradigm that enables collaborators to be partner-aware—capable of adapting to specific intervention agents through selective incorporation of helpful suggestions while maintaining invariance to misleading ones—thereby developing the "intentionality" necessary for robust collaborative reasoning.
+[2023] terms “intentionality,” or the capacity to autonomously evaluate interventions based on their causal impact on task outcomes rather than superficial plausibility.
+To address this limitation, we need a learning paradigm that enables collaborators to be partner-aware, or capable of adapting to specific intervention agents through selective incorporation of helpful suggestions while maintaining invariance to misleading ones—thereby developing the "intentionality" necessary for robust collaborative reasoning.
 Such a collaborator would maintain reasoned agency in the face of various intervention qualities, leading to more robust collaboration and better common ground convergence across diverse interaction scenarios.
-In other words, effective collaborators must remain safely interruptible [Orseau and Armstrong, 2016]a delicate balance between receptive and robust that renders them open to incorporating valuable insights that genuinely contribute to task success, yet capable of maintaining their reasoning integrity when faced with misleading suggestions.
+In other words, effective collaborators must remain safely interruptible [Orseau and Armstrong, 2016].
+This is a delicate balance between receptive and robust that renders them open to incorporating valuable insights that genuinely contribute to task success, yet capable of maintaining their reasoning integrity when faced with misleading suggestions.
 This motivates our Interruptible Collaborative Roleplayer (ICR) learning algorithm.
@@
 Theoretically, as $\lambda_{\text{Intent}}\rightarrow\infty$, $\Delta_{\text{CF}}(\pi_{C}^{CI})$ approaches zero, making our policy’s performance approach that of the optimal policy $\pi_{C}^{*}$ (subject to task optimization constraints $\epsilon_{task}$).
-This theoretical guarantee connects directly to Lemma A.2 showing that while preference-aligned policies achieve Bellman optimality at both token and intervention levels in the underlying MDP, they remain suboptimal in the MAMDP due to failing to account for intervention quality.
+This theoretical guarantee connects directly to Lemma˜B.2 showing that while preference-aligned policies achieve Bellman optimality at both token and intervention levels in the underlying MDP, they remain suboptimal in the MAMDP due to failing to account for intervention quality.
 Our counterfactual invariance objective $\mathcal{J^{*}}(\theta_{C})$ bridges this gap by explicitly teaching collaborator LLM agents to distinguish between interventions based on their causal impact on task outcomes during training, rather than merely using their in-context learning capacity.
@@
 2 and 5 for prompts.
-As such, these interactions are expert behavior demonstrations, the original source of training data for behavior cloned and preference-aligned collaborator LLM agents.
+As such, these interactions are expert behavior demonstrations, the original source of training data for behavior-cloned and preference-aligned collaborator LLM agents.
 For evaluation, all trained ICR collaborators and competing baselines are first deployed following the MAMDP interaction in the expert data collection, and then evaluated, primarily on their ability to reach consensus during the collaboration.
@@
 Therefore, to control for language interpretability, rather than using full natural language, collaborator agents act over a discrete space of structured beliefs—allowing us to evaluate grounded reasoning without requiring fluency.
-In the Weights Task, collaborators express beliefs as symbolic propositions over block weights (i.e., green > red, blue = 10g), while in DeliData, agents select from predefined stances—support, oppose, unsure, or consider_latertoward questions of which cards to flip.
+In the Weights Task, collaborators express beliefs as symbolic propositions over block weights (i.e., green > red, blue = 10g), while in DeliData, agents select from predefined stances toward questions of which cards to flip: support, oppose, unsure, or consider_later.
 Agents are trained independently using only task-specific proxy rewards: factual accuracy in the Weights domain, and logically aligned card-checking in DeliData (e.g., +1 for supporting parts of a correct solution—a vowel or odd number, -1 for incorrect support, +0.5 for justified uncertainty).
@@
 For Weights Task (WTD), we report accuracy scores (ACC), a composite metric that multiplies the percentage of correct propositions by the total size of the common ground, rewarding both factual accuracy and common-ground convergence, while penalizing trivial solutions.
-For example, ACC of 14 indicates that the collaborator agents were able to recover 14 out a 37 total theoretically possible propositions at the end of the collaboration, adjusted for correctness.
+For example, ACC of 14 indicates that the collaborator agents were able to recover 14 out of a 37 total theoretically possible propositions at the end of the collaboration, adjusted for correctness.
 For DeliData, we report both accuracy (ACC)—a task-specific fine-grained score [Karadzhov et al., 2024] based on the final submission (after $N=15$ turns)—and common ground gain (CG), defined as the net increase in unique solution types introduced during the dialogue beyond those initially proposed.
@@
 Agent Baseline ACC ACC ACC CG ACC CG
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
+BC-Collaborator 5.97±0.05 6.04±0.07 0.71±0.02 -0.13±0.18 0.68±0.03 -0.15±0.19
+DPO 9.56±0.09 7.60±0.09 0.82±0.02 2.80±0.19 0.79±0.02 2.65±0.20
+IPO 7.64±0.07 6.80±0.07 0.78±0.02 2.87±0.21 0.75±0.02 2.72±0.22
+PPO 7.37±0.09 7.81±0.11 0.81±0.02 2.94±0.18 0.78±0.03 2.80±0.19
+PSO-Intent 8.09±0.08 6.35±0.09 0.76±0.03 2.73±0.20 0.73±0.03 2.58±0.21
+ICR 14.06±0.13 10.87±0.13 0.88±0.02 3.35±0.19 0.85±0.02 3.18±0.20
 Table 1: Performance across collaborator-agent baselines interacting with a fixed intervention agent over 100 dialogues (15 turns each) on DeliData and Weights Task (WTD).
@@
 Any remaining errors are the responsibility of the authors.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4.
+Justification: The abstract and introduction clearly state the paper’s main contributions i.e., modeling collaborator–intervention agent dynamics via the MAMDP framework and introducing the ICR training method with counterfactual KL regularization.
+These are supported throughout the methodology and experiments across both tasks.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9.
+Justification: Yes, we provide a separate limitations (integrated with future work ) section in sec.
+7 that mentions limitations in training and evaluation of LLM agents in our settings that is integrated with future work directions.
+We also mention that we could only evaluate on two task domains with 8b scale LLMs due to compute budgets.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [Yes]
+  14.
+Justification: Yes, we provide theoretical results with detailed proofs and assumptions in Theorem˜B.4, Theorem˜3.2, Lemma˜3.1 and Lemma˜B.2 (kept it in the appendix since its not the core result).
+We tried to keep them abridged in the main paper but detailed assumptions and statements are provided in the appendix.
+Appendix˜B.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19.
+Justification: We describe the experimental section including dataset and task-benchmarks in details in  Section˜5 and Appendix˜D including all prompts used for experimentation.
+Code to run the experiment will be provided in supplementary material.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24.
+Justification: We describe the experimental section including dataset and task-specific evaluation metrics in details in  Section˜5 and training related hyperparameters in Appendix˜D including all prompts used for experimentation.
+Code to run the experiment will be provided in supplementary material.
+Expert data collection using LLMs and prompts and diversity settings like temperature and top-p for reproduction of results is also provided in Appendix˜D.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: All training related hyperparameters including LLM architecture, Lora settings, optimizer, batch size, iterations, are provided in Appendix˜D
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34.
+Justification: Table˜1 and Figure˜1(a) provides is our main result and we report the standard error over 100 collaboration trials, each with 15 turns.
+Figure˜1(b) provides ICR proxy “training” reward evolution $\lambda_{\text{intent}}$ values in ablating across 3 runs (with 3 different random seeds total).
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: Refer Appendix˜D for these details.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: We have reviewed it and confirm that our paper is consistent with it.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49.
+Justification: We provide positive social impacts in the introduction, motivation especially with respect to how collaborative agents can help learning in multi-party collaborations.
+We do not clearly see any immediate negative impacts of our work.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [No]
+  54.
+Justification: We do not explicitly provide safeguards since we are mostly testing a proof-of-concept idea, but ideas in safe-interruptibility that we engage are positively correlated with safety-related aspects of LLM usage.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: Yes, we cite the LLMs used in the paper in our experiments section section˜5 as well as the original source of the dataset used.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [Yes]
+  64.
+Justification: Our data-generation method is well-documented in our experiments section (section˜5) and original datasets used for sourcing bootstrap dialogues (to prompt the experts) are cited in various sections in the paper.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: We do not conduct any crowdsourcing or human evaluations.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: No human experiments are conducted in our work.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer:[N/A]
+  79. Justification: We used LLMs primarily for formatting, occasional paraphrasing and grammar refinement, and, in some cases, for assisting with data visualization.
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
 Appendix A Additional Results Under Alternative Evaluation Conditions
@@
 Agent Baseline ACC ACC CG
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
+ICR-Masked (with GPT-4o) 7.23±0.11 0.75±0.04 2.15±0.31
+ICR-Small (with Llama 3-8B-Instruct) 8.45±0.13 0.80±0.03 2.45±0.30
+PSO-Skeptical (with GPT-4o) 6.89±0.10 0.74±0.03 2.01±0.27
+ICR-Phrasing (with GPT-4o) 12.34±0.17 0.84±0.03 3.08±0.28
+PPO-CF (with GPT-4o) 8.34±0.16 0.79±0.03 2.56±0.31
+GPT-4o-mini (with GPT-4o-mini) 12.47±0.21 0.79±0.03 2.78±0.25
+GPT-4o (with GPT-4o) 15.23±0.21 0.91±0.03 3.34±0.25
 Table 2: Performance across alternative baselines and evaluation conditions on 50 DeliData and Weights Task dialogues (15 turns each) in the full-press setting only. Format follows Table 1.
@@
 These results suggest that, first, having a strong intervention agent like GPT-4o leads to optimal ICR performance across both tasks.
-However, ICR agents are still capable of leveraging weaker intervention agents compared to no interventions at all, as shown by the improvement from masked interventions (ICR-Masked: $7.23\textsubscript{$\pm$0.11}$ Weights, $75\%$ DeliData accuracy) to weak intervention agents (ICR-Small: $8.45\textsubscript{$\pm$0.13}$ Weights, $80\%$ DeliData accuracy).
-Second, the PSO-Skeptical baseline shows a slight degradation in performance across both tasks ($6.89\textsubscript{$\pm$0.10}$ Weights, $74\%$ DeliData with $2.01\textsubscript{$\pm$0.27}$ common ground) when using negative polarity prompting, compared to standard PSO with positive polarity prefix (see PSO-Intent in Table 1), according to Ward et al.
+However, ICR agents are still capable of leveraging weaker intervention agents compared to no interventions at all, as shown by the improvement from masked interventions (ICR-Masked: $7.23\textsubscript{±0.11}$ Weights, $75\%$ DeliData accuracy) to weak intervention agents (ICR-Small: $8.45\textsubscript{±0.13}$ Weights, $80\%$ DeliData accuracy).
+Second, the PSO-Skeptical baseline shows a slight degradation in performance across both tasks ($6.89\textsubscript{±0.10}$ Weights, $74\%$ DeliData with $2.01\textsubscript{±0.27}$ common ground) when using negative polarity prompting, compared to standard PSO with positive polarity prefix (see PSO-Intent in Table 1), according to Ward et al.
 [2023]’s strategy.
@@
 These results suggest that ICR’s improved performance is an effect of the objective rather than the extra training.
-Third, ICR-Phrasing ($12.34\textsubscript{$\pm$0.17}$ Weights, $84\%$ DeliData with $3.08\textsubscript{$\pm$0.28}$ common ground)—which simply swaps out the counterfactual prefix—demonstrates ICR’s robustness to semantic variations in counterfactual phrasing across both collaborative reasoning tasks.
+Third, ICR-Phrasing ($12.34\textsubscript{±0.17}$ Weights, $84\%$ DeliData with $3.08\textsubscript{±0.28}$ common ground)—which simply swaps out the counterfactual prefix—demonstrates ICR’s robustness to semantic variations in counterfactual phrasing across both collaborative reasoning tasks.
 The variance from the main results under different wordings is low, and at no point does ICR’s performance dip within the margin of error of any other method reported in Table 1.
-Additionally, standard PPO with a simple counterfactual prompt addition (PPO-CF) achieves $8.34\textsubscript{$\pm$0.16}$ Weights, $79\%$ DeliData, and lags behind ICR training since ICR explicitly makes agents robust to counterfactual framing via policy gradient methods.
+Additionally, standard PPO with a simple counterfactual prompt addition (PPO-CF) achieves $8.34\textsubscript{±0.16}$ Weights, $79\%$ DeliData, and lags behind ICR training since ICR explicitly makes agents robust to counterfactual framing via policy gradient methods.
 Using standard PPO with simple prompt augmentation can confuse the model, since the model is forced to pay attention to both standard as well as counterfactually-based contexts, without specific counterfactual regularization.
 This could explain the performance drop in this case, whereas ICR’s counterfactual regularization term mitigates this effect.
-Finally, expert agents like GPT-4o achieve strong performance when paired together ($15.23\textsubscript{$\pm$0.21}$ in Weights task and $91\%$ accuracy in DeliData with $3.34\textsubscript{$\pm$0.25}$ common ground), though this may reflect GPT-4o’s extensive pretraining on reasoning tasks, potentially including exposure to DeliData or DeliData-like problems.
+Finally, expert agents like GPT-4o achieve strong performance when paired together ($15.23\textsubscript{±0.21}$ in Weights task and $91\%$ accuracy in DeliData with $3.34\textsubscript{±0.25}$ common ground), though this may reflect GPT-4o’s extensive pretraining on reasoning tasks, potentially including exposure to DeliData or DeliData-like problems.
 Our human evaluation of GPT-4o in these tasks (Section˜D.4) shows high agreement with humans, supporting our choice of this expert model.
-The GPT-4o-mini pairing shows competitive performance ($12.47\textsubscript{$\pm$0.21}$ Weights, $79\%$ DeliData) compared to standard baselines—though lower than ICR as well as the larger GPT-4o model—demonstrating that expert model collaboration can achieve strong results across both collaborative reasoning domains.
+The GPT-4o-mini pairing shows competitive performance ($12.47\textsubscript{±0.21}$ Weights, $79\%$ DeliData) compared to standard baselines—though lower than ICR as well as the larger GPT-4o model—demonstrating that expert model collaboration can achieve strong results across both collaborative reasoning domains.
 It is important to note that ICR in our main experiments uses Llama 3-8B-Instruct, and so we can see that a weaker base model trained with ICR performs comparably with GPT-4o, even including GPT-4o’s potential prior exposure to the task, and the implicit advantage that comes with using GPT-4o as both intervener and collaborator.
@@
 Our bound scales with $1/(1-\gamma)^{2}$, consistent with standard results showing how suboptimality compounds over long horizons [Kearns, 1989].
-The analysis incorporates a causal influence parameter $C$ that quantifies how strongly interventions affect optimal task behavior, similar to the influence measures developed in trust-based [Fung et al., 2024] and causal [Jaques et al., 2019].
+The analysis incorporates a causal influence parameter $C$ that quantifies how strongly interventions affect optimal task behavior, similar to the influence measures developed in trust-based [Fung et al., 2024] and causal [Jaques et al., 2019] methods.
 Appendix C Prompts